### v1

# Total variation distance between SDEs with stable noise and Brownian motion with applications to Poisson PDEs

Abstract.
We consider a $d$-dimensional stochastic differential equation (SDE) of the form $dU_{t}=b(U_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}Z_{t}$ with initial condition $U_{0}=u\in{\mathds{R}^{d}}$ and additive noise. Denote by $X_{t}^{x}$ the (unique) solution with initial value $X_{0}^{x}=x$, if the driving noise $Z_{t}$ is a $d$-dimensional rotationally symmetric $\alpha$-stable ($21<\alpha<$ 2) Lévy process, and by $Y_{t}^{y}$ the (unique) solution with initial value $Y_{0}^{y}=y$, if the driving noise is a $d$-dimensional Brownian motion.
We give precise estimates of the total variation distance $\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\|_{\mathrm{TV}}$ for all $t>0$, and we show that the ergodic measures $\mu_{\alpha}$ and $\mu_{2}$ of $X_{t}^{x}$ and $Y_{t}^{y}$, respectively, satisfy $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right]$.
As an application, we consider the following $d$-dimensional Poisson PDE
$-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
where $h\in C_{b}({\mathds{R}^{d}},\real)$ and $1<\alpha\leq 2$. We show that this PDE admits a solution $f_{\alpha}$, which is continuous, but may have linear growth. For fixed $\alpha_{0}\in(1,2)$, we show that for any $\alpha_{0}\leq\alpha<2$ there is a constant $C(\alpha_{0})$ such that
$\left|f_{\alpha}(x)-f_{2}(x)\right|\leq C(\alpha_{0})(1+|x|)\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha}\quad\text{for all $x$}.$
1. Introduction
This paper can be seen as a continuation of our previous paper [5], where we studied the 1-Wasserstein distance between stochastic differential equations (SDEs) with different drivers based on the methods developed in [1]. This means that we shall work in the same set-up and under similar assumptions, which we will briefly recall.
1.1. General setup and notation
Consider the SDEs
(1.1) $\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}L_{t},\quad X_{0}=x,$
(1.2) $\mathrm{d}Y_{t}=b(Y_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}B_{t},\quad Y_{0}=y,$
driven by a standard $d$-dimensioal Brownian motion $B_{t}$ and a stable Lévy process $L_{t}$ with characteristic function $\mathds{E}\,\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, $21<\alpha<$ 2; this is a standard rotationally symmetric $\alpha$-stable process run at half speed. The drift coefficient $b:{\mathds{R}^{d}}\to{\mathds{R}^{d}}$ will be specified further down, $\sigma$ is an invertible $d\times d$ matrix.
Our main assumptions are as follows:
(H1).
There exist constants $\theta_{0}>0$ and $K\geq 0$ such that
$\langle x-y,b(x)-b(y)\rangle\leq-\theta_{0}|x-y|^{2}+K\quad\text{for all\ \ }x,y\in{\mathds{R}^{d}};$
(H2).
$b\in C^{2}({\mathds{R}^{d}},{\mathds{R}^{d}})$ and there exist constants $\theta_{1},\theta_{2}\geq 0$ such that for all $x,v,w\in{\mathds{R}^{d}}$,
(1.3) $|\nabla_{v}b(x)|$ $\leq\theta_{1}|v|,$
(1.4) $|\nabla_{w}\nabla_{v}b(x)|$ $\leq\theta_{2}|v||w|.$
If $f:{\mathds{R}^{d}}\to\real$ is smooth enough, the directional derivative in direction $v\in{\mathds{R}^{d}}$ is defined as
$\nabla_{v}f(x):=\lim_{\epsilon\to 0}\frac{f(x+\epsilon v)-f(x)}{\epsilon}=\nabla f(x)\cdot v$
and $\nabla_{w}\nabla_{v}f(x):=\nabla_{w}(\nabla_{v}f)(x)=\nabla^{2}f(x)v\cdot w$ for $v,w\in{\mathds{R}^{d}}$. The vector and matrix norms of $\nabla f$ and $\nabla^{2}f$ are given by
$|\nabla f(x)|=\sup_{v\in{\mathds{R}^{d}},|v|=1}\left|\nabla_{v}f(x)\right|\quad\text{and}\quad\|\nabla f(x)\|_{\mathrm{op}}=\sup_{v,w\in{\mathds{R}^{d}},|v|=|w|=1}\left|\nabla_{w}\nabla_{v}f(x)\right|$
[…]
$\nabla f(x)v$ $=\nabla_{v}f(x)=(\nabla_{v}f_{1}(x),\dots,\nabla_{v}f_{d}(x))^{\top},$
$\nabla^{2}f(x)vu$ $=\nabla_{u}\nabla_{v}f(x)=(\nabla_{u}\nabla_{v}f_{1}(x),\dots,\nabla_{u}\nabla_{v}f_{d}(x))^{\top}.$
The Hilbert–Schmidt inner product of two matrices $A,B\in{}^{d\times d}$ is $\left\langle A,B\right\rangle_{\mathrm{HS}}=\sum_{i,j=1}^{d}A_{ij}B_{ij}$, and the Hilbert-Schmidt norm is $\|A\|_{\mathrm{HS}}=\sqrt{\sum_{i,j=1}^{d}A_{ij}^{2}}$.
Finally, $\mathrm{Lip}(\lambda)$ is the family of Lipschitz functions $f:{\mathds{R}^{d}}\to\real$ with Lipschitz constant $\lambda$ and norm $\|f\|_{\mathrm{Lip}}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|}\leq\lambda$.
We will frequently need the constants
(1.5) $A(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{d/2}\Gamma(1-\frac{\alpha}{2})},\quad\omega_{d-1}:=\frac{2\pi^{d/2}}{\Gamma\left(\frac{d}{2}\right)},$
which can be estimated as follows (see [5, Lemma 3.3]):
(1.6) $\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}\leq C\quad\text{and}\quad\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\leq C(2-\alpha)\log(1+d).$
Under (1.3) the both SDEs (1.1) and (1.2) have unique non-explosive (strong) solutions, which we denote by $X_{t}$ and $Y_{t}$, respectively. These are time homogeneous Markov processes and write $X_{t}^{x}$ and $Y_{t}^{y}$ to indicate the initial position $X_{0}^{x}=x$ and $Y_{0}^{y}=y$. The operator semigroup induced by the Markov process $(X_{t}^{x})_{t\geq 0}$ is
$P_{t}f(x)=\mathds{E}f(X^{x}_{t}),\quad f\in C_{b}({\mathds{R}^{d}},\real),\,t>0.$
We need to study $P_{t}$ on the Banach space $\left(C_{\mathrm{lin}}({\mathds{R}^{d}},\real),\|\cdot\|_{\mathrm{lin}}\right)$ of linearly growing continuous functions
$C_{\mathrm{lin}}({\mathds{R}^{d}},\real)=\left\{f\in C({\mathds{R}^{d}},\real)\,;\,\|f\|_{\mathrm{lin}}:=\sup_{x\in{\mathds{R}^{d}}}\frac{|f(x)|}{1+|x|}<\infty\right\}.$
By [5], the semigroup $(P_{t})_{t\geq 0}$ is well defined on $C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$. The infinitesimal generator $\mathscr{A}^{P}$ of the semigroup $(P_{t})_{t\geq 0}$ has the following domain:
$\mathcal{D}(\mathscr{A}^{P}):=\bigg{\{}f\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)\,;\,$ $g(x)=\lim_{t\to 0}\frac{P_{t}f(x)-f(x)}{t}\text{\ \ exists \ forall $x\in{\mathds{R}^{d}}$}$
$\text{and\ }g\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)\bigg{\}}.$
In the same way we can consider the semigroup $Q_{t}$ of $(Y_{t}^{y})_{t\geq 0}$ and its generator $\left(\mathscr{A}^{Q},\mathcal{D}(\mathscr{A}^{Q})\right)$.
From the classical Lyapunov function criterion, see e.g. [10], we know that under (H1) and (H2) the solutions to the SDEs (1.1) and (1.2) are ergodic. Denote by $\mu_{\alpha}$ and $\mu_{2}$ the respective ergodic measures of $X_{t}^{x}$ and $Y_{t}^{y}$.
1.2. Main results
We will now describe the principal results of our paper. Write $\mathscr{C}_{b}(1)=\{h\in C_{b}({\mathds{R}^{d}},\real)\>;\>\|h\|_{\infty}\leq 1\}\}$ for the unit ball in $C_{b}({\mathds{R}^{d}},\real)$. The total variation distance of two probability measures (on ${\mathds{R}^{d}}$) $\mu$ and $\nu$ is given by
$\|\mu-\nu\|_{\mathrm{TV}}:=\sup_{h\in\mathscr{C}_{b}(1)}|\mu(h)-\nu(h)|$
(as usual we write $\mu(h):=\int h\,\mathrm{d}\mu$ etc.). It is clear that $\|\mu-\nu\|_{\mathrm{TV}}\leq 2$ for all probability measures $\mu$ and $\nu$.
Throughout this paper, $C,C_{1},C_{2}$ denote positive constants which may depend on $\theta_{0},\theta_{1},\theta_{2},K,\|\sigma\|_{\mathrm{op}},\|\sigma^{-1}\|_{\mathrm{op}}$, but they are always independent of $d$ and $\alpha$; their values may change from line to line.
Theorem 1.1.
Assume (H1) and (H2). For any $\alpha\in(1,2)$, $x,y\in{\mathds{R}^{d}}$ and $t>0$,
$\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\|_{\mathrm{TV}}$
$\qquad\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|+\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
In particular,
(1.7) $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right].$
In the estimate (1.7) we cannot hope, in general, for a faster speed of convergence than $2-\alpha$. This can be seen from the one-dimensional Ornstein–Uhlenbeck, cf. Proposition 7.1 below.
We can apply this theorem to study the convergence rate of Poisson PDEs as follows ($1<\alpha\leq 2$):
(1.8) $-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
where $h\in C_{b}({\mathds{R}^{d}},\real)$,
$-(-\Delta)^{\alpha/2}f(x)=\int_{{\mathds{R}^{d}}\setminus\{0\}}\left[f(x+z)-f(x)-\langle\nabla f(x),z\rangle\mathds{1}_{(0,1)}(|z|)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
is the fractional Laplacian for $\alpha\in(1,2)$, and $\Delta f(x)=\sum_{i=1}^{d}\partial^{2}_{i}f(x)$ is the usual Laplace operator. This type of Poisson equation has many applications in central limit theorems related to SDEs and approximations of probability measures, see [9] and [7, 20].
Our second theorem is as follows:
Theorem 1.2.
Assume (H1) and (H2).
  1. a)
    For any $\alpha\in(1,2]$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the equation (1.8) admits a solution $f_{\alpha}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$.
  2. b)
    Let $\alpha_{0}\in(1,2)$ be arbitrary. For any $\alpha\in[\alpha_{0},2)$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the solution $f_{\alpha}$ from a) satisfies
    $\|f_{\alpha}-f_{2}\|_{\mathrm{lin}}\leq C(\alpha_{0})\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha},$
    with a constant $C(\alpha_{0})$ depending on $\alpha_{0},\theta_{0},\theta_{1},\theta_{2},K$.
2. The key gradient estimate for Theorem 1.1
This section is devoted to proving the following gradient estimate, which is crucial for the proof of our main result.
Proposition 2.1.
Assume (H2). For any $t\in(0,1]$, $\alpha\in(0,2)$ and $f\in C^{2}_{b}({\mathds{R}^{d}},\real)$,
[…]
2.1. A random time-change method
This method was originally proposed by X. Zhang [22] and, in a similar context used in our publication [1]. In order to keep the presentation self-contained, we recall the basic ingredients here. Using Bochner’s subordination, we can in law represent the stable Lévy process $(L_{t})_{t\geq 0}$ with characteristic function $\exp(-\frac{1}{2}t|\xi|^{\alpha})$ as a time-changed Brownian motion $(W_{S_{t}})_{t\geq 0}$, where $(W_{t})_{t\geq 0}$ is a standard Brownian motion with characteristic function $\exp(-\frac{1}{2}t|\xi|^{2})$ and $(S_{t})_{t\geq 0}$ is an independent stable subordinator (i.e. an +-valued Lévy process) withthe following Laplace transform:
$\mathds{E}\left[\mathrm{e}^{-rS_{t}}\right]=\mathrm{e}^{-2^{-1}t(2r)^{\alpha/2}},\quad r>0,\,t\geq 0,$
see e.g. [2, 15, 22]. Since we are interested only in distributional properties of the solution to the SDE (1.1), we may replace it by
(2.1) $\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{S_{t}},\quad X_{0}=x.$
This allows us to work with the canonical realizations of Brownian motion and the subordinator on the product probability space $\left(\mathds{W}\times\mathds{S},\mathscr{B}(\mathds{W})\otimes\mathscr{B}(\mathds{S}),\mu_{\mathds{W}}\times\mu_{\mathds{S}}\right)$, where $\mathds{W}$ and $\mathds{S}$ are the canonical path spaces of Brownian motion: $\mathds{W}$ are the continuous functions $w:[0,\infty)\to{\mathds{R}^{d}}$ with $w(0)=0$ and $\mathds{S}$ are the right-continuous increasing functions $\ell:[0,\infty)\to[0,\infty)$ with $\ell(0)=0$. In particular, we can calculate the expectation of the solution $X_{t}^{x}$ to the SDE (2.1) as
(2.2) $\mathds{E}f\left(X_{t}^{x}\right)=\int_{\mathds{S}}\int_{\mathds{W}}f\left(X_{t}^{x}(w\circ\ell)\right)\mu_{\mathds{W}}(\mathrm{d}w)\,\mu_{\mathds{S}}(\mathrm{d}\ell).$
This means that we can freeze the path of the subordinator and do all calculations for a Brownian motion which is time-changed with a deterministic time-change $\ell=(\ell_{t})_{t\geq 0}\in\mathds{S}$. After that, we only need to make sure that our results remain stable under the integration with respect to $\mu_{\mathds{S}}(\mathrm{d}\ell)$. If needed, we write $\mathds{E}^{\mathds{W}}$ and $\mathds{E}^{\mathds{S}}$ for the expectation taken only in the $\mathds{W}$ or $\mathds{S}$ component.
For fixed $\ell\in\mathds{S}$, denote by $X_{t}^{\ell;x}$ the solution to the SDE
(2.3) $\mathrm{d}X_{t}^{x;\ell}=b(X_{t}^{x;\ell})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{\ell_{t}-\ell_{0}},\quad X_{0}^{x;\ell}=x.$
By construction, $W_{\ell_{t}}$ is a càdlàg $\mathscr{F}_{\ell_{t}}^{\mathds{W}}$-martingale. Thus, (H1) and (H2) guarantee that the SDE (2.3) admits a unique càdlàg $\mathscr{F}_{\ell_{t}}^{\mathds{W}}$-adapted solution $X_{t}^{x;\ell}$ for each initial value $x\in{\mathds{R}^{d}}$.
For technical reasons, we have to regularize $\ell\in\mathds{S}$. Fix $\epsilon\in(0,1)$ and define
$\ell_{t}^{\epsilon}:=\frac{1}{\epsilon}\int_{t}^{t+\epsilon}\ell_{s}\,\mathrm{d}s+\epsilon t=\int_{0}^{1}\ell_{\epsilon s+t}\,\mathrm{d}s+\epsilon t.$
It is not hard to see that $\ell_{t}^{\epsilon}\downarrow\ell_{t}$ ($t\geq 0$) as $\epsilon\downarrow 0$ and $t\mapsto\ell_{t}^{\epsilon}$ is absolutely continuous and strictly increasing. In slight abuse of notation we write $X_{t}^{x;\ell^{\epsilon}}$ for the solution to the SDE (2.3) with $\ell$ replaced by $\ell^{\epsilon}$.
Let $\gamma^{\epsilon}$ be the inverse function of $\ell^{\epsilon}$, i.e., $\ell^{\epsilon}\circ\gamma^{\epsilon}_{t}=t$ for all $t\geq\ell_{0}^{\epsilon}$ and $\gamma^{\epsilon}\circ\ell^{\epsilon}_{t}=t$ for all $t\geq 0$. By construction, $t\mapsto\gamma_{t}^{\epsilon}$ is absolutely continuous on $[\ell_{0}^{\epsilon},\infty)$. Define
$Z_{t}^{x;\ell^{\epsilon}}:=X_{\gamma_{t}^{\varepsilon}}^{x;\ell^{\epsilon}},\quad t\geq\ell_{0}^{\epsilon}.$
From (2.3) (with $\ell\rightsquigarrow\ell^{\epsilon}$) and with a change of variables, we see for $t\geq\ell_{0}^{\epsilon}$,
(2.4) $Z_{t}^{x;\ell^{\epsilon}}=x+\int_{0}^{\gamma_{t}^{\epsilon}}b\left(X_{s}^{x;\ell^{\epsilon}}\right)\mathrm{d}s+\sigma W_{t-\ell^{\epsilon}_{0}}=x+\int_{\ell_{0}^{\epsilon}}^{t}b\left(Z_{s}^{x;\ell^{\epsilon}}\right)\dot{\gamma}_{s}^{\epsilon}\,\mathrm{d}s+\sigma W_{t-\ell^{\epsilon}_{0}}.$
2.2. Integration by parts via Malliavin Calculus
For fixed $t>\ell^{\epsilon}_{0}$, consider the Hilbert space $H:=L^{2}([\ell^{\epsilon}_{0},t];{\mathds{R}^{d}})$ equipped with the inner product
$\langle f,g\rangle_{H}:=\int_{\ell^{\epsilon}_{0}}^{t}\langle f(s),g(s)\rangle\,\mathrm{d}s\quad\text{for $f,g\in H$}.$
Let $W=\{W(h)\,;\,h\in H\}$ be an isonormal Gaussian process associated with $H$, i.e.
$\mathds{E}^{\mathds{W}}[W(f)W(g)]=\langle f,g\rangle_{H},\quad f,g\in H,$
which is defined on the Wiener space $\left(\mathds{W},\mathscr{B}(\mathds{W}),\mu_{\mathds{W}}\right)$. For the $i$th standard coordinate vector $e_{i}\in{\mathds{R}^{d}}$, $i=1,2,\dots,d$, it is clear that $W(e_{1}\mathds{1}_{[\ell^{\epsilon}_{0},t]}(s)),\dots,W(e_{d}\mathds{1}_{[\ell^{\epsilon}_{0},t]}(s))$ are pairwise independent one-dimensional standard Brownian motions.
Let $D$ denote the Malliavin derivative; the directional Malliavin derivative of a random variable $F\in\mathds{D}^{1,2}$ (the domain of $D$ in $L^{2}(\mathds{W},\mathscr{B}(\mathds{W}),\mu_{\mathds{W}})$) along $u\in H$ is
$D_{u}F:=\langle DF,u\rangle_{H}=\int_{\ell^{\epsilon}_{0}}^{t}\langle DF(s),u(s)\rangle\,\mathrm{d}s$
with $DF\in L^{2}(\mathds{W},H)$. For a $d$-dimensional random vector $F=(F_{1},\dots,F_{d})^{\top}$ with $F_{i}\in\mathds{D}^{1,2}$ for each $1\leq i\leq d$, we set $DF:=(DF_{1},\dots DF_{d})^{\top}$ and $D_{u}F=\left(D_{u}F_{1},\dots,D_{u}F_{d}\right)^{\top}$.
Let $\delta$ be the divergence operator, i.e. the adjoint of $D$. It is well known that, for any $u\in\mathrm{Dom}(\delta)$ and $F\in\mathds{D}^{1,2}$, $\delta(u)$ satisfies
(2.5) $\mathds{E}^{\mathds{W}}\left[F\delta(u)\right]=\mathds{E}^{\mathds{W}}[\langle DF,u\rangle_{H}].$
The following auxiliary lemma is crucial for proving Proposition 2.1.
Lemma 2.2.
Assume (H2), and fix $\ell\in\mathds{S}$ and $\epsilon\in(0,1)$. For any $t>0$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell^{\epsilon}}_{t}\right)\right]\right|\leq C\left(\frac{1}{\sqrt{\ell^{\epsilon}_{t}-\ell^{\epsilon}_{0}}}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell^{\epsilon}}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
Proof.
By (2.4), the Malliavin derivative of $Z^{x;\ell^{\epsilon}}_{t}$ along the direction $u\in H$ satisfies
$\mathrm{d}D_{u}Z^{x;\ell^{\epsilon}}_{t}=\left[\nabla b\left(Z^{x;\ell^{\epsilon}}_{t}\right)\dot{\gamma}^{\epsilon}_{t}D_{u}Z^{x;\ell^{\epsilon}}_{t}+\sigma u(t)\right]\,\mathrm{d}t,\quad t\geq\ell^{\epsilon}_{0}.$
Since $D_{u}Z^{x;\ell^{\epsilon}}_{\ell^{\epsilon}_{0}}=0$, this equation has a solution of the form
(2.6) $D_{u}Z^{x;\ell^{\epsilon}}_{t}=\int_{\ell^{\epsilon}_{0}}^{t}J_{s,t}\sigma u(s)\,\mathrm{d}s,$
where
$J_{s,t}:=\exp\left(\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right),\quad\ell^{\epsilon}_{0}\leq s\leq t.$
For any fixed $t>\ell^{\epsilon}_{0}$ and $1\leq i\leq d$, let
(2.7) $u_{i}(s):=\frac{1}{t-\ell^{\epsilon}_{0}}\,\sigma^{-1}J_{s,t}^{-1}e_{i}.$
It can be easily verified that
$D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}=e_{i},\quad 1\leq i\leq d.$
Combining this with the duality (2.5) between the Malliavin derivative and the divergence operator gives us, for any $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $1\leq i\leq d$,
$\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]$ $=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),e_{i}\big{\rangle}\right]=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}\big{\rangle}\right]$
$=\mathds{E}^{\mathds{W}}\left[D_{u_{i}}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]=\mathds{E}^{\mathds{W}}\left[f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\delta\left(u_{i}\right)\right].$
By the Cauchy–Schwarz inequality and [12, Proposition 1.3.1], we obtain
(2.8) $\begin{split}\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|&\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\delta(u_{i})^{2}\right]}\\ &\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\|u_{i}\|_{H}^{2}+\|Du_{i}\|_{H\otimes H}^{2}\right]},\end{split}$
where $\|\cdot\|_{H\otimes H}$ denotes the norm on the tensor space $H\otimes H$. By (1.3) and (2.7), one has
(2.9) $\begin{split}\|u_{i}\|_{H}^{2}&=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\exp\left(-\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)e_{i}\right|^{2}\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\exp\left(2\theta_{1}\int_{s}^{t}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{t-\ell^{\epsilon}_{0}}\,\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}.\end{split}$
Similarly, using (1.3) and (2.6) gives for any $v\in H$
(2.10) $\left|D_{v}Z^{x;\ell^{\epsilon}}_{t}\right|\leq\|\sigma\|_{\mathrm{op}}\|v\|_{H}\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{t-\ell^{\epsilon}_{0}}.$
Furthermore, by the chain rule for matrix exponential functions (see e.g. [13]), one has
$\begin{split}D_{v}J_{s,t}^{-1}&=\int_{0}^{1}\exp\left(-\tau\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\left(-\int_{s}^{t}\nabla^{2}b\left(Z^{x;\ell^{\epsilon}}_{r}\right)D_{v}Z^{x;\ell^{\epsilon}}_{r}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\mbox{}\times\exp\left(-(1-\tau)\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}\tau,\end{split}$
which, together with (1.3), (1.4) and (2.10), yields that
$\|D_{v}J_{s,t}^{-1}\|_{\mathrm{op},\infty}\leq\|\sigma\|_{\mathrm{op}}\theta_{2}\|v\|_{H}\sqrt{t-\ell^{\epsilon}_{0}}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}},\quad s\in[\ell^{\epsilon}_{0},t].$
Then, it follows from the chain rule and (2.7) that
$\|D_{v}u_{i}\|^{2}_{H}$ $=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}}^{t}\left|\sigma^{-1}\left\{D_{v}J_{s,t}^{-1}\right\}e_{i}\right|^{2}\mathrm{d}s$
$\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}\|v\|_{H}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
Since this holds for any $v\in H$, one can deduce that
$\|Du_{i}\|^{2}_{H\otimes H}\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
Combining this with (2.8) and (2.9) gives us for $t\geq\ell^{\epsilon}_{0}$,
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|$
$\quad\leq\left(\frac{\|\sigma^{-1}\|_{\mathrm{op}}}{\sqrt{t-\ell^{\epsilon}_{0}}}\,\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}+\|\sigma^{-1}\|_{\mathrm{op}}\|\sigma\|_{\mathrm{op}}\theta_{2}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}$
$\quad\leq C\left(\frac{1}{\sqrt{t-\ell^{\epsilon}_{0}}}+\gamma^{\epsilon}_{t}\right)\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}.$
Since $Z_{\ell_{t}^{\epsilon}}^{x;\ell^{\epsilon}}=X_{t}^{x;\ell^{\epsilon}}$, we can replace $t$ by $\ell_{t}^{\epsilon}$ and get the desired estimate. ∎
2.3. Proof of Proposition 2.1
Lemma 2.3.
Assume (H2). For any $\alpha\in(0,2)$, $t\in(0,1]$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
[…]
Proof.
Letting $\epsilon\downarrow 0$ in Lemma 2.2 and using $\ell^{\epsilon}_{t}\downarrow\ell_{t}$ and [22, Lemma 2.2], we get for $t>0$,
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\right|\leq C\left(\ell_{t}^{-1/2}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
Now we “unfreeze” the fixed path $\ell$ and use (2.2). From the Cauchy–Schwarz inequality we get for $t>0$,
$\left|P_{t}(\partial_{i}f)(x)\right|$ $=\left|\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)\right|$
$\leq\int_{\mathds{S}}\big{|}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\big{|}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C\mathrm{e}^{2\theta_{1}t}\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)^{2}\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)}$
$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1}+t^{2}\right)\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\mathds{E}\left[\big{|}f(X^{x}_{t})\big{|}^{2}\right]}$
$=C\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}\left[S_{t}^{-1}\right]+t^{2}}\,\sqrt{P_{t}|f|^{2}(x)}$
$\leq C\mathrm{e}^{2\theta_{1}t}\left(\sqrt{\mathds{E}\left[S_{t}^{-1}\right]}+t\right)\sqrt{P_{t}|f|^{2}(x)}.$
From [4, Lemma 4.1] we know the following moment identity for $\alpha\in(0,2)$
(2.11) $\mathds{E}\left[S_{t}^{-1}\right]=\frac{1}{2}\,\Gamma\left(1+\tfrac{2}{\alpha}\right)2^{2/\alpha}t^{-2/\alpha},\quad t>0.$
Substituting this identity into the previous inequality, we obtain the claimed estimate for $t\in(0,1]$. ∎
Proof of Proposition 2.1.
It follows from Lemma 2.3 that for all $t\in(0,1]$, $\alpha\in(0,2)$ and $x\in{\mathds{R}^{d}}$,
$\left\|\mathds{E}\left[\nabla^{2}f(X^{x}_{t})\right]\right\|_{\mathrm{HS}}^{2}$ $=\sum_{i,j=1}^{d}\left|\mathds{E}\left[\partial_{i}\partial_{j}f(X^{x}_{t})\right]\right|^{2}$
[…]
Finally,
$\sum_{j=1}^{d}P_{t}|\partial_{j}f|^{2}(x)=P_{t}|\nabla f|^{2}(x)\leq\|\nabla f\|_{\infty}^{2},$
and this completes the proof. ∎
3. Gradient estimates of $Q_{t}$ and comparison of the generators
In this section we collect further tools for the proof of Theorem 1.1. Recall that $\mathscr{C}_{b}(1)$ is the unit ball in $C_{b}({\mathds{R}^{d}},\real)$.
Lemma 3.1.
Assume (H1) and (H2). For all $h\in\mathscr{C}_{b}(1)$ and $t>0$,
(3.1) $\|\nabla Q_{t}h\|_{\infty}$ $\leq C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t},$
(3.2) $\|\nabla^{2}Q_{t}h\|_{\mathrm{op},\infty}$ $\leq C_{1}\left\{t^{-1}\vee 1\right\}\mathrm{e}^{-C_{2}t},$
where both $C_{1}$ and $C_{2}$ do not depend on the dimension $d$.
Proof.
According to [6, Corollary 2], for any $t>0$ and $x,y\in{\mathds{R}^{d}}$,
$\sup_{f\in\mathrm{Lip}(1)}\left|Q_{t}f(x)-Q_{t}f(y)\right|\leq C_{1}\mathrm{e}^{-C_{2}t}|x-y|,$
which means that for any $f\in\mathrm{Lip}(1)$,
$\|\nabla Q_{t}f\|_{\infty}\leq C_{1}\mathrm{e}^{-C_{2}t}\quad\text{for all $t>0$}.$
On the other hand, it follows from [14] that for all $h\in\mathscr{C}_{b}(1)$ and $t\in(0,1]$,
$\|\nabla Q_{t}h\|_{\infty}\leq Ct^{-1/2},$
which implies that $Q_{t}h$ is a Lipschitz function with Lipschitz constant $Ct^{-1/2}$. Hence, if $t\in(0,1]$, (3.1) holds. If $t>1$, we have
$\|\nabla Q_{t}h\|_{\infty}=\|\nabla Q_{t-1}\{Q_{1}h\}\|_{\infty}\leq CC_{1}\mathrm{e}^{-C_{2}(t-1)}=CC_{1}\mathrm{e}^{C_{2}}\mathrm{e}^{-C_{2}t},$
which yields (3.1).
To prove (3.2), we first note that by [5, Lemma 2.1, (2.2)], for any $f\in\mathrm{Lip}(1)$ and $t\in(0,1]$,
$\|\nabla^{2}Q_{t}f\|_{\mathrm{op},\infty}\leq Ct^{-1/2}.$
So by (3.1), if $t\in(0,2]$, it holds that for $h\in\mathscr{C}_{b}(1)$
$\|\nabla^{2}Q_{t}h\|_{\mathrm{op},\infty}=\|\nabla^{2}Q_{t/2}\{Q_{t/2}h\}\|_{\mathrm{op},\infty}$ $\leq C_{1}(t/2)^{-1/2}\mathrm{e}^{-C_{2}t/2}C(t/2)^{-1/2}$
$=2C_{1}Ct^{-1}\mathrm{e}^{-C_{2}t/2},$
while for $t>2$ we have for $h\in\mathscr{C}_{b}(1)$
$\|\nabla^{2}Q_{t}h\|_{\mathrm{op},\infty}=\|\nabla^{2}Q_{1}\{Q_{t-1}h\}\|_{\mathrm{op},\infty}\leq C_{1}\mathrm{e}^{-C_{2}(t-1)}C=C_{1}C\mathrm{e}^{C_{2}}\mathrm{e}^{-C_{2}t}.$
This completes the proof of (3.2). ∎
Lemma 3.2.
Assume (H1) and (H2).
  1. a)
    For any $\alpha\in(1,2)$ and $0<s<t\leq 1$,
    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}$
    $\qquad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d.$
  2. b)
    For any $\alpha\in(1,2)$, $0<s<t\leq 1$ and $k=1,2,\dots$,
    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\leq C_{1}d^{3/2}(2-\alpha)\left((t-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{e}^{-C_{2}k}.$
Proof.
We begin with some general preparations. Let $h\in\mathscr{C}_{b}(1)$ and set $f:=Q_{s+k}h$ for $k\in\{0,1,2,\dots\}$. It is not hard to see that the generators of the semigroups $(P_{t})_{t\geq 0}$ and $(Q_{t})_{t\geq 0}$ are given by
$\mathscr{A}^{P}f(x)=\langle\nabla f(x),b(x)\rangle+\int\limits_{\mathclap{{\mathds{R}^{d}}\setminus\{0\}}}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\mathds{1}_{(0,1)}(|z|)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z,$
and
$\mathscr{A}^{Q}f(x)=\langle\nabla f(x),b(x)\rangle+\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}.$
Therefore,
$\left(\mathscr{A}^{P}-\mathscr{A}^{Q}\right)f(x)$
=$=\left\{\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
$\quad\mbox{}+\int_{|z|\geq 1}\left[f(x+\sigma z)-f(x)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
$=:\mathsf{J}_{1}(x)+\mathsf{J}_{2}(x).$
By the first inequality in (1.6), we get for all $\alpha\in(1,2)$,
(3.3) $\begin{split}|\mathsf{J}_{2}(x)|&\leq\int_{|z|\geq 1}\left|f(x+\sigma z)-f(x)\right|\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq\|\nabla f\|_{\infty}\|\sigma\|_{\mathrm{op}}\int_{|z|\geq 1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\|\nabla f\|_{\infty}\|\sigma\|_{\mathrm{op}}\frac{A(d,\alpha)\omega_{d-1}}{\alpha-1}\\ &\leq\frac{C}{\alpha-1}\,\|\nabla f\|_{\infty}(2-\alpha)d.\end{split}$
We rewrite $\mathsf{J}_{1}(x)$ in the following form:
$\mathsf{J}_{1}(x)$
$=\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}$
$=\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x+r\sigma z),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}$
=$=\left\{\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
$\qquad\mbox{}+\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x+r\sigma z)-\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
$=:\mathsf{J}_{11}(x)+\mathsf{J}_{12}(x).$
Using the symmetry of the measure $\rho(\mathrm{d}z)=|z|^{-d-\alpha}\mathrm{d}z$, it is clear that $\int_{|z|\leq 1}z_{i}z_{j}\,\rho(\mathrm{d}z)=\delta_{ij}\frac{1}{d}\int_{|z|\leq 1}|z|^{2}\,\rho(\mathrm{d}z)$, and so we get
$\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
$\qquad=\frac{1}{2}\int_{|z|<1}\langle\nabla^{2}f(x),\sigma(zz^{\top})\sigma^{\top}\rangle_{\mathrm{HS}}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
$\qquad=\frac{1}{2}\,\frac{1}{d}\int_{|z|<1}\langle\nabla^{2}f(x),|z|^{2}\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
$\qquad=\frac{1}{2d}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\int_{|z|<1}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$\qquad=\frac{A(d,\alpha)\omega_{d-1}}{2d(2-\alpha)}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}.$
Combining this with (1.6), Proposition 2.1 and the fact that $\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\leq\sqrt{d}\|\sigma\sigma^{\top}\|_{\mathrm{op}}\leq\sqrt{d}\|\sigma\|_{\mathrm{op}}^{2}$, we obtain for all $\alpha\in(1,2)$,
(3.4) $\begin{split}|P_{t-s}\mathsf{J}_{11}(x)|&=\left|\frac{1}{2}\left[\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right]\big{\langle}P_{t-s}\nabla^{2}f(x),\sigma\sigma^{\top}\big{\rangle}_{\mathrm{HS}}\right|\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{HS}}\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\\ &\leq C(2-\alpha)(t-s)^{-1/\alpha}\|\nabla f\|_{\infty}d\log(1+d).\end{split}$
Let us now estimate $P_{t-s}\mathsf{J}_{12}(x)$. Since $\|(\sigma z)(\sigma z)^{\top}\|_{\mathrm{HS}}=|\sigma z|^{2}$, it follows from Proposition 2.1 that for any $r\in[0,1]$ and $\alpha\in(1,2)$,
$\left|\left\langle P_{t-s}\left[\nabla^{2}\left(f(x+r\sigma z)-f(x)\right)\right],(\sigma z)(\sigma z)^{\top}\right\rangle_{\mathrm{HS}}\right|$
$\quad\leq\left\|P_{t-s}\nabla^{2}\{f(\cdot+r\sigma z)-f(\cdot)\}(x)\right\|_{\mathrm{HS}}\left\|(\sigma z)(\sigma z)^{\top}\right\|_{\mathrm{HS}}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\|\nabla\{f(\cdot+r\sigma z)-f(\cdot)\}\right\|_{\infty}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{[2\|\nabla f\|_{\infty}]\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}r|\sigma z|]\right\}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\right\}|z|^{2}.$
This yields for $\alpha\in(1,2)$,
(3.5) $\begin{split}&|P_{t-s}\mathsf{J}_{12}(x)|\\ &\leq\int_{|z|<1}\int_{0}^{1}\left|\langle P_{t-s}\nabla^{2}f(x+r\sigma z)-P_{t-s}\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}\right|(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left(\int_{0}^{1}(1-r)\,\mathrm{d}r\right)\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\frac{C}{2}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z.\end{split}$
Now we are ready to prove Lemma 3.2.
Proof of 3.2.a): Take in the above calculations $k=0$. Note that $f=Q_{s}h$ and $s\in(0,1)$. By (3.1), we get $\|\nabla f\|_{\infty}\leq Cs^{-1/2}$. Then by (3.3) and (3.4), we have for all $\alpha\in(1,2)$,
$|\mathsf{J}_{2}(x)|\leq\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
and
$|P_{t-s}\mathsf{J}_{11}(x)|\leq C(2-\alpha)(t-s)^{-1/\alpha}s^{-1/2}d\log(1+d),$
Moreover, by (3.2), $\|\nabla^{2}f\|_{\mathrm{op},\infty}\leq Cs^{-1}$. It follows from the first inequality in (1.6) that for all $\alpha\in(0,2)$,
$\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}\leq Cd.$
Therefore, we obtain from (3.5) that for $\alpha\in(1,2)$,
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left\{s^{-1/2}\wedge[s^{-1}|z|]\right\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C(t-s)^{-1/\alpha}\sqrt{d}\left(\int_{|z|\leq\sqrt{s}}s^{-1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z+\int_{\sqrt{s}<|z|<1}s^{-1/2}\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}\sqrt{d}\,A(d,\alpha)\omega_{d-1}\left(s^{-1/2}\int_{0}^{\sqrt{s}}r^{2-\alpha}\,\mathrm{d}r+\int_{\sqrt{s}}^{1}r^{1-\alpha}\,\mathrm{d}r\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\sqrt{d}\,\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
$\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}.$
Combining all estimates, we get for all $\alpha\in(1,2)$, $x\in{\mathds{R}^{d}}$ and $0<s<t\leq 1$,
$\left|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})f(x)\right|$
$\quad\leq|P_{t-s}\mathsf{J}_{11}(x)|+|P_{t-s}\mathsf{J}_{12}(x)|+|P_{t-s}\mathsf{J}_{2}(x)|$
$\quad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left[(2-\alpha)+\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\right]d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
from which we immediately get the first part of the lemma.
Proof of 3.2.b): Let $k\geq 1$. Since $f=Q_{s+k}h$ and $s+k>1$, it follows from Lemma 3.1 that
$\|\nabla f\|_{\infty}\leq C_{1}\mathrm{e}^{-C_{2}k}\quad\text{and}\quad\|\nabla^{2}f\|_{\mathrm{op},\infty}\leq C_{1}\mathrm{e}^{-C_{2}k}.$
Using (3.3) and (3.4) implies for all $\alpha\in(1,2)$,
$|\mathsf{J}_{2}(x)|\leq\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
and
$|P_{t-s}\mathsf{J}_{11}(x)|\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d\log(1+d).$
With the first estimate in (1.6), it is easy to check that for any $\alpha\in(0,2)$ the following estimate holds:
$\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}\leq C(2-\alpha)d.$
Using (3.5), we get for all $\alpha\in(1,2)$,
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C_{1}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|\leq 1}\mathrm{e}^{-C_{2}k}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}A(d,\alpha)\omega_{d-1}\int_{0}^{1}r^{2-\alpha}\,\mathrm{d}r$
$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}$
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}.$
Finally, combining all estimates, we conclude that for all $\alpha\in(1,2)$,
$\left|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})f(x)\right|$ $\leq|P_{t-s}\mathsf{J}_{11}(x)|+|P_{t-s}\mathsf{J}_{12}(x)|+|P_{t-s}\mathsf{J}_{2}(x)|$
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}+\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
which completes the proof of the second part of the lemma. ∎
4. Proof of Theorem 1.1
Lemma 4.1.
For any $t\in(0,1]$ and $\alpha\in(1,2)$,
$\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\leq\frac{C}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).$
Proof.
Using the elementary estimate
$1-s^{\kappa}=-\log s\int_{0}^{\kappa}s^{x}\,\mathrm{d}x\leq(-\log s)\kappa=\kappa\log\tfrac{1}{s},\qquad 0<s<1,\;\kappa>0,$
yields for all $s\in(0,1)$ and $\alpha\in(0,2)$,
$3-\alpha-\frac{s^{1-\frac{\alpha}{2}}}{3-\alpha}$ $=\frac{(4-\alpha)(2-\alpha)}{3-\alpha}+\frac{1}{3-\alpha}\left(1-s^{1-\frac{\alpha}{2}}\right)$
$<2(2-\alpha)+\left(1-\tfrac{\alpha}{2}\right)\log\tfrac{1}{s}$
$<2(2-\alpha)\left(1+\log\tfrac{1}{s}\right).$
This, together with the change of variables $s=t\tau$, implies that for any $t\in(0,1]$ and $\alpha\in(1,2)$,
(4.1) $\begin{split}\int_{0}^{t}&(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\leq 2(2-\alpha)\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(1+\log\tfrac{1}{s}\right)\mathrm{d}s\\ &=2(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\int_{0}^{1}(1-\tau)^{-\frac{1}{\alpha}}\tau^{-\frac{1}{2}}\left(1+\log\tfrac{1}{t\tau}\right)\mathrm{d}\tau\\ &=2(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(\int_{0}^{1}(1-\tau)^{-\frac{1}{\alpha}}\tau^{-\frac{1}{2}}\left(1+\log\tfrac{1}{\tau}\right)\mathrm{d}\tau+\operatorname{B}\left(1-\tfrac{1}{\alpha},\tfrac{1}{2}\right)\log\tfrac{1}{t}\right).\end{split}$
As usual, $\operatorname{B}(\cdot,\cdot)$ is Euler’s Beta function. From $\Gamma(x)<1/x$ for $x\in(0,1)$, we see that for all $\alpha\in(1,2)$,
(4.2) $\operatorname{B}\left(1-\tfrac{1}{\alpha},\tfrac{1}{2}\right)=\frac{\Gamma\left(1-\frac{1}{\alpha}\right)\Gamma\left(\frac{1}{2}\right)}{\Gamma\left(\frac{3}{2}-\frac{1}{\alpha}\right)}<\sqrt{\pi}\,\Gamma\left(1-\tfrac{1}{\alpha}\right)<\sqrt{\pi}\,\frac{\alpha}{\alpha-1}<\frac{2\sqrt{\pi}}{\alpha-1}.$
Noting that there exists some constant $C>0$ such that
$1+\log\tfrac{1}{\tau}\leq C\tau^{-1/4}\quad\text{for all $\tau\in(0,1)$},$
as in (4.2) we have for all $\alpha\in(1,2)$,
$\int_{0}^{1}(1-\tau)^{-\frac{1}{\alpha}}$ $\tau^{-\frac{1}{2}}\left(1+\log\tfrac{1}{\tau}\right)\mathrm{d}\tau\leq C\int_{0}^{1}(1-\tau)^{-\frac{1}{\alpha}}\tau^{-\frac{3}{4}}\,\mathrm{d}\tau$
$=C\operatorname{B}\left(1-\tfrac{1}{\alpha},\tfrac{1}{4}\right)=C\frac{\Gamma\left(1-\frac{1}{\alpha}\right)\Gamma\left(\frac{1}{4}\right)}{\Gamma\left(\frac{5}{4}-\frac{1}{\alpha}\right)}$
$<C\frac{\Gamma\left(\frac{1}{4}\right)}{\Gamma\left(\frac{3}{4}\right)}\,\Gamma\left(1-\tfrac{1}{\alpha}\right)<\frac{\Gamma\left(\frac{1}{4}\right)}{\Gamma\left(\frac{3}{4}\right)}\,\frac{2C}{\alpha-1}.$
Substituting this and (4.2) into (4.1), finishes the proof. ∎
Proposition 4.2.
Assume (H1) and (H2). For all $\alpha\in(1,2)$ and $t>0$,
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
Proof.
Recall that $\mathscr{C}_{b}(1)$ is the unit ball in $C_{b}({\mathds{R}^{d}},\real)$. Since
$\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}|P_{t}h(x)-Q_{t}h(x)|$
and
$P_{t}h-Q_{t}h=-\int_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}s}\,P_{t-s}Q_{s}h\,\mathrm{d}s=\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\,\mathrm{d}s,$
we get
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\sup_{x\in{}^{d}}\left|\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h(x)\,\mathrm{d}s\right|.$
From Lemmas 3.2 and 4.1 we see that for all $\alpha\in(1,2)$ and $t\in(0,1]$,
(4.3) $\begin{split}\sup_{h\in\mathscr{C}_{b}(1)}&\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\leq Cd^{3/2}\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\quad\mbox{}+\frac{Cd}{\alpha-1}(2-\alpha)\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\\ &\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right)+\frac{2Cd}{\alpha-1}(2-\alpha)t^{1/2}\\ &\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).\end{split}$
Combining this with
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s,$
proves assertion for $t\in(0,1]$.
Next, we consider the case $t>1$. Denote by $\lfloor t\rfloor$ the largest integer which is less than or equal to $t>1$. From
$\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\,\mathrm{d}s$
$=\sum_{k=0}^{\lfloor t\rfloor-1}\int_{k}^{k+1}P_{t-k-1}P_{k+1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\,\mathrm{d}s+\int_{\lfloor t\rfloor}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\,\mathrm{d}s$
$=\sum_{k=0}^{\lfloor t\rfloor-1}\int_{0}^{1}P_{t-k-1}P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\,\mathrm{d}s+\int_{0}^{t-\lfloor t\rfloor}P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\,\mathrm{d}s,$
we obtain
(4.4) $\begin{split}\sup_{x\in{}^{d}}&\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\\ &\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sum_{k=1}^{\lfloor t\rfloor-1}\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\right\|_{\infty}\mathrm{d}s,\end{split}$
where we use the ‘empty-sum-convention’ $\sum_{k=1}^{0}\dots=0$. Using (4.3) with $t=1$, we see that for all $\alpha\in(1,2)$,
$\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
By Lemma 3.2.b), we find for all $\alpha\in(1,2)$,
$\sup_{h\in\mathscr{C}_{b}(1)}$ $\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s$
$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\int_{0}^{1}\left((1-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k},$
and
$\sup_{h\in\mathscr{C}_{b}(1)}$ $\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+\lfloor t\rfloor}h\right\|_{\infty}\mathrm{d}s$
$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\int_{0}^{t-\lfloor t\rfloor}\left((t-\lfloor t\rfloor-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\left((t-\lfloor t\rfloor)^{1-1/\alpha}\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}(t-\lfloor t\rfloor)\right)$
$\leq C_{1}d^{3/2}(2-\alpha)\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha).$
Substituting these estimates into (4.4), shows that for all $t\geq 1$ and $\alpha\in(1,2)$,
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}$ $\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\left(1+\sum_{k=1}^{\lfloor t\rfloor-1}\mathrm{e}^{-C_{2}k}\right)$
$\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
This completes the proof. ∎
Proposition 4.3.
Assume (H1) and (H2). For all $x,y\in{\mathds{R}^{d}}$ and $t>0$,
$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|\right].$
Proof.
By the definition of the total variation norm $\|\cdot\|_{\mathrm{TV}}$, one has
$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\left|Q_{t}h(x)-Q_{t}h(y)\right|\leq\sup_{h\in\mathscr{C}_{b}(1)}\left\|\nabla Q_{t}h\right\|_{\infty}|x-y|.$
Together with (3.1) we get the claimed estimate. ∎
Proof of Theorem 1.1.
Note that
$\left\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\right\|_{\mathrm{TV}}\leq\left\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\right\|_{\mathrm{TV}}+\left\|\mathrm{law}(Y_{t}^{x})-\mathrm{law}(Y_{t}^{y})\right\|_{\mathrm{TV}}.$
Therefore, the first assertion follows straight from Propositions 4.2 and 4.3.
From the classical ergodic theory for Markov processes (see, for instance, [18]), one has
$\lim_{t\to\infty}\left\|\mathrm{law}(X_{t}^{x})-\mu_{\alpha}\right\|_{\mathrm{TV}}=\lim_{t\to\infty}\left\|\mathrm{law}(Y_{t}^{y})-\mu_{2}\right\|_{\mathrm{TV}}=0,\quad x,y\in{\mathds{R}^{d}}.$
Since for any fixed $x,y\in{\mathds{R}^{d}}$
$\left\|\mu_{\alpha}-\mu_{2}\right\|_{\mathrm{TV}}\leq\left\|\mu_{\alpha}-\mathrm{law}(X_{t}^{x})\right\|_{\mathrm{TV}}+\left\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\right\|_{\mathrm{TV}}+\left\|\mathrm{law}(Y_{t}^{y})-\mu_{2}\right\|_{\mathrm{TV}},$
we can let $t\to\infty$ to get (1.7) from the first part. ∎
5. Moment estimates for Theorem 1.2
The following two moment estimates for the solutions to the SDEs (1.2) and (1.1) are needed to obtain moment estimates for the ergodic measures. Note that under our assumptions on the drift coefficient, the solution $Y_{t}^{y}$ to the SDE (1.2) has finite second moments, but we will need a precise estimate of the second moment.
Lemma 5.1.
Assume (H1). The following moment estimate holds for the solution $(Y_{t}^{y})_{t\geq 0}$ to the SDE (1.2):
$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq\mathrm{e}^{-\theta_{0}t}|y|^{2}+Cd\left(1-\mathrm{e}^{-\theta_{0}t}\right),\quad\text{for all\ \ }y\in{\mathds{R}^{d}},\;t\geq 0.$
Proof.
From (H1) we get for any $x\in{\mathds{R}^{d}}$,
(5.1) $\begin{split}\langle b(x),x\rangle&\leq-\theta_{0}|x|^{2}+\langle b(0),x\rangle+K\\ &\leq-\theta_{0}|x|^{2}+\sqrt{\theta_{0}}|x|\frac{|b(0)|}{\sqrt{\theta_{0}}}+K\\ &\leq-\theta_{0}|x|^{2}+\frac{\theta_{0}}{2}|x|^{2}+\frac{|b(0)|^{2}}{2\theta_{0}}+K\\ &=-\frac{\theta_{0}}{2}|x|^{2}+C_{1},\end{split}$
where $C_{1}:=|b(0)|^{2}/(2\theta_{0})+K$. By Itô’s formula,
$|Y_{t}^{y}|^{2}$ $=|y|^{2}+\int_{0}^{t}\left(2\langle Y_{s}^{y},b(Y_{s}^{y})\rangle+\|\sigma\sigma^{\top}\|_{\mathrm{HS}}^{2}\right)\mathrm{d}s+2\int_{0}^{t}Y_{s}^{y}\,\mathrm{d}B_{s}$
$\leq|y|^{2}+\int_{0}^{t}\left(-\theta_{0}|Y_{s}^{y}|^{2}+2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d\right)\mathrm{d}s+2\int_{0}^{t}Y_{s}^{y}\,\mathrm{d}B_{s}.$
Using the finiteness of the second moment we get that
$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]=|y|^{2}+\int_{0}^{t}\mathds{E}\left[2\langle Y_{s}^{y},b(Y_{s}^{y})\rangle+\|\sigma\sigma^{\top}\|_{\mathrm{HS}}^{2}\right]\mathrm{d}s;$
in particular $\mathds{E}\left[|Y_{t}^{y}|^{2}\right]$ is differentiable in $t$. Thus, the above inequality shows
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq-\theta_{0}\mathds{E}\left[|Y_{t}^{y}|^{2}\right]+\left(2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d\right).$
By the Gronwall inequality, we see that for all $t>0$,
$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq\mathrm{e}^{-\theta_{0}t}|y|^{2}+\frac{2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d}{\theta_{0}}\left(1-\mathrm{e}^{-\theta_{0}t}\right).$
This gives the desired estimate. ∎
Lemma 5.2.
Assume (H1). The following moment estimate holds for the solution $(X_{t}^{x})_{t\geq 0}$ to the SDE (1.1) with $\alpha\in(1,2)$:
$\mathds{E}\left[\left(1+|X_{t}^{x}|^{2}\right)^{1/2}\right]\leq\mathrm{e}^{-\theta_{0}t/2}\left(1+|x|^{2}\right)^{1/2}+\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right)\quad\text{for all\ \ }x\in{\mathds{R}^{d}},\;t\geq 0.$
Proof.
Set $f(\cdot):=\sqrt{1+|\cdot|^{2}}$. By Dynkin’s formula,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}f(X_{t}^{x})=\mathds{E}\mathscr{A}^{P}f(X_{t}^{x})=\mathds{E}\left[\mathsf{J}_{1}+\mathsf{J}_{2}+\mathsf{J}_{3}\right],$
where
$\mathsf{J}_{1}$ $:=\langle\nabla f(X_{t}^{x}),b(X_{t}^{x})\rangle,$
$\mathsf{J}_{2}$ $:=\int_{|z|\geq 1}\left[f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z,$
$\mathsf{J}_{3}$ $:=\int_{|z|<1}\left[f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})-\langle\nabla f(X_{t}^{x}),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z.$
We will estimate the terms $\mathsf{J}_{i}$, $i=1,2,3$, separately. From (5.1) we get
$\mathsf{J}_{1}$ $=\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}\langle X_{t}^{x},b(X_{t}^{x})\rangle$
$\leq\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}\left(-\frac{\theta_{0}}{2}|X_{t}^{x}|^{2}+C_{1}\right)$
$=-\frac{\theta_{0}}{2}\left(1+|X_{t}^{x}|^{2}\right)^{1/2}+\left(C_{1}+\frac{\theta_{0}}{2}\right)\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}$
$\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+C_{2},$
where $C_{2}:=C_{1}+\theta_{0}/2$. Since
$\langle 2X_{t}^{x}+\sigma z,\sigma z\rangle\leq(|X_{t}^{x}+\sigma z|+|X_{t}^{x}|)|\sigma z|\leq\left(\sqrt{1+|X_{t}^{x}+\sigma z|^{2}}+\sqrt{1+|X_{t}^{x}|^{2}}\right)|\sigma z|,$
we have
$f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})=\frac{\langle 2X_{t}^{x}+\sigma z,\sigma z\rangle}{\sqrt{1+|X_{t}^{x}+\sigma z|^{2}}+\sqrt{1+|X_{t}^{x}|^{2}}}\leq|\sigma z|.$
Therefore,
$\mathsf{J}_{2}\leq\|\sigma\|_{\mathrm{op}}\int_{|z|\geq 1}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z=\|\sigma\|_{\mathrm{op}}\frac{A(d,\alpha)\omega_{d-1}}{\alpha-1}.$
A straightforward calculation shows for all $\xi,\eta\in{\mathds{R}^{d}}$,
$\left\langle\nabla^{2}f(\xi),\eta\eta^{\top}\right\rangle_{\mathrm{HS}}$ $=(1+|\xi|^{2})^{-3/2}\sum_{i,j=1}^{d}\left[(1+|\xi|^{2})\delta_{ij}-\xi_{i}\xi_{j}\right]\eta_{i}\eta_{j}$
$=(1+|\xi|^{2})^{-3/2}\left(\sum_{i=1}^{d}\left[1+|\xi|^{2}-\xi_{i}^{2}\right]\eta_{i}^{2}-\sum_{i=1}^{d}\xi_{i}\eta_{i}\sum_{j\neq i}\xi_{j}\eta_{j}\right)$
$\leq(1+|\xi|^{2})^{-3/2}\left((1+|\xi|^{2})\sum_{i=1}^{d}\eta_{i}^{2}+\left(\sum_{i=1}^{d}|\xi_{i}\eta_{i}|\right)^{2}\right)$
$\leq(1+|\xi|^{2})^{-3/2}\left((1+|\xi|^{2})|\eta|^{2}+|\xi|^{2}|\eta|^{2}\right)$
$\leq 2(1+|\xi|^{2})^{-1/2}|\eta|^{2}$
$\leq 2|\eta|^{2}.$
Thus, we get for all $z\in{\mathds{R}^{d}}$,
$f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})-\langle\nabla f(X_{t}^{x}),\sigma z\rangle$ $=\int_{0}^{1}\left\langle\nabla^{2}f(X_{t}^{x}+r\sigma z),(\sigma z)(\sigma z)^{\top}\right\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r$
$\leq 2|\sigma z|^{2}\int_{0}^{1}(1-r)\,\mathrm{d}r$
$=|\sigma z|^{2}.$
This implies that
$\mathsf{J}_{3}\leq\|\sigma\|_{\mathrm{op}}^{2}\int_{|z|<1}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z=\|\sigma\|_{\mathrm{op}}^{2}\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}.$
Now we combine the estimates for the terms $\mathsf{J}_{i}$, $i=1,2,3$, and apply (1.6). Just as in the proof of Lemma 5.1 (where the second moment is now replaced by $\mathds{E}f(X_{t}^{x})$) we conclude that for all $\alpha\in(1,2)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}f(X_{t}^{x})$ $\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+C_{2}+\|\sigma\|_{\mathrm{op}}\frac{A(d,\alpha)\omega_{d-1}}{\alpha-1}+\|\sigma\|_{\mathrm{op}}^{2}\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
$\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+\frac{Cd}{\alpha-1}.$
Gronwall’s inequality finally yields
$\mathds{E}f(X_{t}^{x})\leq\mathrm{e}^{-\theta_{0}t/2}f(x)+\frac{2}{\theta_{0}}\,\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right).\qed$
Proposition 5.3.
Assume (H1). Then
$\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)\leq C\sqrt{d}\quad\text{and}\quad\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)\leq\frac{Cd}{\alpha-1}\text{\ \ for all $\alpha\in(1,2)$}.$
Proof.
By the classical ergodic theory for Markov processes, see e.g. [11], we have for all $x,y\in{\mathds{R}^{d}}$,
$\lim_{t\to\infty}W_{1}\left(\mathrm{law}(X_{t}^{x}),\mu_{\alpha}\right)=\lim_{t\to\infty}W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)=0,$
where $W_{1}$ is the Wasserstein-$1$ distance, which is for two probability measures $\mu,\nu$ given by
$W_{1}(\mu,\nu)=\sup_{f\in\mathrm{Lip}(1)}|\mu(f)-\nu(f)|.$
By the triangle inequality and Lemma 5.1, we obtain for all $y\in{\mathds{R}^{d}}$ and $t\geq 0$,
$\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)$ $\leq\left|\mathds{E}\left[|Y_{t}^{y}|\right]-\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)\right|+\mathds{E}\left[|Y_{t}^{y}|\right]$
$\leq W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)+\sqrt{\mathds{E}\left[|Y_{t}^{y}|^{2}\right]}$
$\leq W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)+\mathrm{e}^{-\theta_{0}t/2}|y|+C\sqrt{d\left(1-\mathrm{e}^{-\theta_{0}t}\right)}.$
Letting $t\to\infty$ we get the first assertion. Similarly, using Lemma 5.2, we have for all $x\in{\mathds{R}^{d}}$, $t\geq 0$ and $\alpha\in(1,2)$,
$\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)$ $\leq\left|\mathds{E}\left[|X_{t}^{x}|\right]-\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)\right|+\mathds{E}\left[|X_{t}^{x}|\right]$
$\leq W_{1}\left(\mathrm{law}(X_{t}^{x}),\mu_{\alpha}\right)+\mathrm{e}^{-\theta_{0}t/2}\left(1+|x|^{2}\right)^{1/2}+\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right).$
Letting $t\to\infty$ finishes the proof. ∎
6. Proof of Theorem 1.2
As before, $C,C_{1},C_{2}$ denote constants whose value may change from line to line. They depend only on the parameters $\theta_{0},\theta_{1},\theta_{2},K$ from (H1) and (H2). If we want to stress further dependences, we write $C(\kappa),C_{1}(\kappa),C_{2}(\kappa)$ etc.
6.1. Gradient estimates for stable semigroup $P_{t}$
Proposition 6.1.
Assume (H1) and (1.3), and let $\sigma=I_{d\times d}$ and $\alpha_{0}\in(1,2)$ be an arbitrary fixed number. For all $\alpha\in[\alpha_{0},2)$, $h\in C({\mathds{R}^{d}},\real)$ and $t>0$, we have
(6.1) $\|\nabla P_{t}h\|_{\infty}\leq C_{1}(\alpha_{0})\left\{t^{-1/\alpha}\vee 1\right\}\mathrm{e}^{-C_{2}(\alpha_{0})t}\|h\|_{\infty}.$
Proof.
Let $S_{t}$ be an $\frac{\alpha}{2}$-stable subordinator with Laplace transform
$\mathds{E}\left[\mathrm{e}^{-rS_{t}}\right]=\mathrm{e}^{-2^{-1}t(2r)^{\alpha/2}},\quad r>0,\,t\geq 0.$
By [3, Corollary 1.2] or [17, Theorem 2.1  (2)], we have for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(0,2)$ and $t>0$,
$\|\nabla P_{t}h\|_{\infty}^{2}\leq\mathds{E}\left[\left(\int_{0}^{t}\mathrm{e}^{-2\theta_{1}r}\,\mathrm{d}S_{r}\right)^{-1}\right]\|h\|_{\infty}\leq\mathrm{e}^{2\theta_{1}t}\mathds{E}\left[S_{t}^{-1}\right]\|h\|_{\infty}.$
Combining this with the moment identity (2.11), we obtain that for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(0,2)$ and $t>0$,
[…]
This means that for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(1,2)$ and $t\in(0,1]$,
(6.2) $\|\nabla P_{t}h\|_{\infty}\leq Ct^{-1/\alpha}\|h\|_{\infty}.$
Thus, (6.1) holds for $t\in(0,1]$ and it remains to prove (6.1) for $t>1$. By Lemma 6.2 and Remark 6.3 below, we have for all $\alpha\in[\alpha_{0},2)$, $x,y\in{}^{d}$ and $t>0$,
$\sup_{f\in\mathrm{Lip}(1)}|P_{t}f(x)-P_{t}f(y)|\leq C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})t}|x-y|,$
which means that for any $f\in\mathrm{Lip}(1)$,
$\|\nabla P_{t}f\|_{\infty}\leq C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})t}\quad\text{for all $t>0$}.$
Since (6.2) implies that $P_{1}h$ is a Lipschitz function with Lipschitz constant $C\|h\|_{\infty}$, we have
$\|\nabla P_{t}h\|_{\infty}=\|\nabla P_{t-1}\{P_{1}h\}\|_{\infty}\leq CC_{1}(\alpha_{0})\|h\|_{\infty}\mathrm{e}^{-C_{2}(\alpha_{0})(t-1)},$
which implies (6.1) for $t>1$. ∎
By a careful calculation following the proof of [19, Theorem 1.2 (i)], we get the following lemma. To save space, we omit the detailed proof here.
Lemma 6.2.
Consider the SDE (1.1) with $\sigma=I_{d\times d}$ and assume that $b:{}^{d}\to{}^{d}$ is a continuous vector field such that for any $x,y\in{}^{d}$,
(6.3) $\langle b(x)-b(y),x-y\rangle\leq\begin{cases}K_{1}|x-y|^{2},&|x-y|\leq L_{0},\\ -K_{2}|x-y|^{2},&|x-y|>L_{0}\end{cases}$
holds with some positive constants $K_{1},K_{2},L_{0}>0$. Let $\alpha_{0}\in(1,2)$ be an arbitrary fixed number. Then there exist constants $C_{i}(\alpha_{0},K_{1},K_{2},L_{0})>0$, $i=1,2$, such that for all $\alpha\in[\alpha_{0},2)$, $x,y\in{}^{d}$ and $t>0$,
$W_{1}\left(\mathrm{law}(X_{t}^{x}),\mathrm{law}(X_{t}^{y})\right)\leq C_{1}(\alpha_{0},K_{1},K_{2},L_{0})\mathrm{e}^{-C_{2}(\alpha_{0},K_{1},K_{2},L_{0})t}|x-y|.$
Remark 6.3.
It is clear that (6.3) follows from (H1) and (1.3); in this case $K_{1}=\theta_{1}$, $K_{2}=\theta_{0}/2$ and $L_{0}=\sqrt{2K/\theta_{0}}$.
6.2. Proof of Theorem 1.2
Proof.
a) By [21, Theorem 1.2 (i)] with $V(x)=1+|x|$ (see also [1, Proposition 1.5]), we know for $\alpha\in(1,2)$,
$\sup_{g:\|g/V\|_{\infty}\leq 1}|P_{t}g(x)-\mu_{\alpha}(g)|\leq C_{1}(d,\alpha)\mathrm{e}^{-C_{2}(d,\alpha)t}V(x),$
which immediately implies that for $h\in C_{b}({\mathds{R}^{d}},\real)$ and $\alpha\in(1,2)$,
(6.4) $|P_{t}h(x)-\mu_{\alpha}(h)|\leq C_{1}(d,\alpha)\mathrm{e}^{-C_{2}(d,\alpha)t}\|h\|_{\infty}(1+|x|).$
It is a well-known result that $Y_{t}$ is ergodic, see e.g. [11], and for $h\in C_{b}({\mathds{R}^{d}},\real)$,
(6.5) $|Q_{t}h(x)-\mu_{2}(h)|\leq C_{1}(d)\mathrm{e}^{-C_{2}(d)t}\|h\|_{\infty}(1+|x|).$
Now we can use a standard method (see e.g. [7, Proposition 6.1]) to show that
(6.6) $f_{\alpha}(x)$ $=\int_{0}^{\infty}\left[P_{t}h(x)-\mu_{\alpha}(h)\right]\mathrm{d}t,$
(6.7) $f_{2}(x)$ $=\int_{0}^{\infty}\left[Q_{t}h(x)-\mu_{2}(h)\right]\mathrm{d}t$
solve (1.8) for $\alpha\in(1,2]$ and $h\in C_{b}({\mathds{R}^{d}},\real)$. It follows from (6.6) and (6.4) that for all $\alpha\in(1,2)$,
$|f_{\alpha}(x)|\leq\int_{0}^{\infty}\left|P_{t}h(x)-\mu_{\alpha}(h)\right|\mathrm{d}t$ $\leq C_{1}(d,\alpha)\|h\|_{\infty}(1+|x|)\int_{0}^{\infty}\mathrm{e}^{-C_{2}(d,\alpha)t}\,\mathrm{d}t$
$=\frac{C_{1}(d,\alpha)}{C_{2}(d,\alpha)}\|h\|_{\infty}(1+|x|),$
hence $f_{\alpha}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$. Similarly, we deduce from (6.7) and (6.5) that $f_{2}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$.
b) Let $h\in C_{b}({\mathds{R}^{d}},\real)$. Without loss of generality, we may assume that $\|h\|_{\infty}\leq 1$, otherwise we replace $h$ by $h/\|h\|_{\infty}$. From (6.6) and (6.7) we have
$|f_{\alpha}(x)-f_{2}(x)|$ $\leq\int_{0}^{\infty}\left|P_{t}h(x)-Q_{t}h(x)-\mu_{\alpha}(h)+\mu_{2}(h)\right|\mathrm{d}t$
$\leq\left|\mu_{\alpha}(h)-\mu_{2}(h)\right|A+\int_{0}^{A}\left|Q_{t}h(x)-P_{t}h(x)\right|\mathrm{d}t$
$\quad\mbox{}+\int_{A}^{\infty}\left|Q_{t}h(x)-\mu_{2}(h)\right|\mathrm{d}t+\int_{A}^{\infty}\left|P_{t}h(x)-\mu_{\alpha}(h)\right|\mathrm{d}t$
$=:\mathsf{I}_{0}+\mathsf{I}_{1}(x)+\mathsf{I}_{2}(x)+\mathsf{I}_{3}(x).$
The number $A\geq 1$ will be chosen later on. We estimate $\mathsf{I}_{0},\dots,\mathsf{I}_{4}(x)$ separately.
From (1.7) we know that for all $\alpha\in[\alpha_{0},2)$,
$\mathsf{I}_{0}\leq\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}A\leq C(\alpha_{0})d^{3/2}(2-\alpha)A.$
For $\mathsf{I}_{1}(x)$ we can use Theorem 1.1 and find for all $\alpha\in[\alpha_{0},2)$ and $t>0$,
$\left|Q_{t}h(x)-P_{t}h(x)\right|$ $\leq\|\mathrm{law}(X^{x}_{t})-\mathrm{law}(Y^{x}_{t})\|_{\mathrm{TV}}$
$\leq C(\alpha_{0})d^{3/2}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
Since for all $\alpha\in[\alpha_{0},2)$,
$\int_{0}^{A}\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\mathrm{d}t\leq\int_{0}^{1}\left(1+t^{-\frac{1}{\alpha_{0}}+\frac{1}{2}}\log\tfrac{1}{t}\right)\mathrm{d}t+\int_{1}^{A}\mathrm{d}t\leq C(\alpha_{0})A,$
we get
$\mathsf{I}_{1}(x)\leq C(\alpha_{0})d^{3/2}(2-\alpha)A,\quad x\in{\mathds{R}^{d}}.$
Because $\mu_{2}$ is an invariant measure of $Q_{t}$, we have for $\mathsf{I}_{2}(x)$
$\int_{A}^{\infty}\left|Q_{t}h(x)-\mu_{2}(h)\right|\mathrm{d}t=\int_{A}^{\infty}\left|\int_{{\mathds{R}^{d}}}\left[Q_{t}h(x)-Q_{t}h(y)\right]\mu_{2}(\mathrm{d}y)\right|\mathrm{d}t.$
For all $t\geq 1$ we have
$\left|Q_{t}h(x)-Q_{t}h(y)\right|$ $=\left|\int_{0}^{1}\langle\nabla Q_{t}h(rx+(1-r)y),x-y\rangle\,\mathrm{d}r\right|$
$\leq\|\nabla Q_{t}h\|_{\infty}|x-y|$
$\leq C_{1}\mathrm{e}^{-C_{2}t}(|x|+|y|),$
where the last inequality is due to (3.1). Hence,
$\mathsf{I}_{2}(x)$ $\leq\int_{A}^{\infty}\int_{\mathds{R}^{d}}|Q_{t}h(x)-Q_{t}h(y)|\,\mu_{2}(\mathrm{d}y)\,\mathrm{d}t$
$\leq C_{1}\int_{\mathds{R}^{d}}(|x|+|y|)\,\mu_{2}(\mathrm{d}y)\int_{A}^{\infty}\mathrm{e}^{-C_{2}t}\,\mathrm{d}t$
$=\frac{C_{1}}{C_{2}}\,\mathrm{e}^{-C_{2}A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{2}(\mathrm{d}y)\right].$
Replacing (3.1) in the above argument with (6.1) yields for all $\alpha\in[\alpha_{0},2)$ the estimate for $\mathsf{I}_{3}(x)$:
$\mathsf{I}_{3}(x)$ $=\int_{A}^{\infty}\left|\int_{\mathds{R}^{d}}[P_{t}h(x)-P_{t}h(y)]\,\mu_{\alpha}(\mathrm{d}y)\right|\mathrm{d}t$
$\leq\int_{A}^{\infty}\int_{\mathds{R}^{d}}|P_{t}h(x)-P_{t}h(y)|\,\mu_{\alpha}(\mathrm{d}y)\,\mathrm{d}t$
$\leq C_{1}(\alpha_{0})\int_{\mathds{R}^{d}}(|x|+|y|)\,\mu_{\alpha}(\mathrm{d}y)\int_{A}^{\infty}\mathrm{e}^{-C_{2}(\alpha_{0})t}\,\mathrm{d}t$
$=\frac{C_{1}(\alpha_{0})}{C_{2}(\alpha_{0})}\,\mathrm{e}^{-C_{2}(\alpha_{0})A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{\alpha}(\mathrm{d}y)\right].$
Altogether we have shown that for all $\alpha\in[\alpha_{0},2)$,
$|f_{\alpha}(x)-f_{2}(x)|$
$\quad\leq C(\alpha_{0})d^{3/2}(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{2}(\mathrm{d}y)+\int_{\mathds{R}^{d}}|y|\,\mu_{\alpha}(\mathrm{d}y)\right]$
$\quad\leq C(\alpha_{0})d^{3/2}(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}d(1+|x|)$
$\quad\leq d^{3/2}\left[C(\alpha_{0})(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}\right](1+|x|),$
where the second inequality is due to Proposition 5.3. Taking $A=1\vee[|\log(2-\alpha)|/C_{2}(\alpha_{0})]$, we get for all $\alpha\in[\alpha_{0},2)$ and $x\in{\mathds{R}^{d}}$,
$|f_{\alpha}(x)-f_{2}(x)|\leq C(\alpha_{0})d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha}\cdot(1+|x|).\qed$
7. Exact rate for the Ornstein–Uhlenbeck process on
In this section, we assume that $\mu_{\alpha}$ is the invariant measure of the Ornstein–Uhlenbeck process on , which solves the following SDE:
$\mathrm{d}X_{t}=-X_{t}\,\mathrm{d}t+\mathrm{d}L_{t},\quad X_{0}=x\in\real.$
Here, $L_{t}$ is a rotationally symmetric $\alpha$-stable Lévy process on with $\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}\right]=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, and $\mu_{2}$ is the invariant measure of
$\mathrm{d}Y_{t}=-Y_{t}\,\mathrm{d}t+\mathrm{d}B_{t},\quad Y_{0}=y\in\real,$
where $B_{t}$ is a standard Brownian motion on with $\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\xi B_{t}}\right]=\mathrm{e}^{-t|\xi|^{2}/2}$.
It is easy to see that (H1) holds with $\theta_{0}=1$ and $K=0$, while (H2) holds with $\theta_{1}=1$ and $\theta_{2}=0$.
Proposition 7.1.
We have
$0<\liminf_{\alpha\uparrow 2}\frac{\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}}{2-\alpha}\leq\limsup_{\alpha\uparrow 2}\frac{\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}}{2-\alpha}<\infty.$
Proof.
The upper bound follows from Theorem 1.1. For the lower bound we use that $X_{t}=\mathrm{e}^{-t}x+\mathrm{e}^{-t}\int_{0}^{t}\mathrm{e}^{s}\,\mathrm{d}L_{s}$. For $\xi\in\real$ we see
$\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\xi X_{t}}\right]$ $=\mathrm{e}^{\mathop{\mathrm{i}}\xi\mathrm{e}^{-t}x}\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\int_{0}^{t}\xi\mathrm{e}^{-t}\mathrm{e}^{s}\,\mathrm{d}L_{s}}\right]=\mathrm{e}^{\mathop{\mathrm{i}}\xi\mathrm{e}^{-t}x}\mathrm{e}^{-2^{-1}\int_{0}^{t}|\xi\mathrm{e}^{-t}\mathrm{e}^{s}|^{\alpha}\,\mathrm{d}s}$
$=\mathrm{e}^{\mathop{\mathrm{i}}\xi\mathrm{e}^{-t}x}\mathrm{e}^{-(2\alpha)^{-1}|\xi|^{\alpha}(1-\mathrm{e}^{-\alpha t})}\;\xrightarrow[]{\;t\to\infty\;}\;\mathrm{e}^{-|\xi|^{\alpha}/(2\alpha)}=\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\xi\alpha^{-1/\alpha}L_{1}}\right].$
Thus, the ergodic measure $\mu_{\alpha}$ is given by the law of $\alpha^{-1/\alpha}L_{1}$. Similarly, $\mu_{2}=\mathrm{law}(2^{-1/2}B_{1})$.
Let
$h(x):=\cos x=\operatorname{Re}\mathrm{e}^{\mathop{\mathrm{i}}x},\quad x\in\real.$
Since $|h(x)|\leq 1$, it follows that
$\|\mu_{2}-\mu_{\alpha}\|_{\mathrm{TV}}$ $\geq\mu_{2}(h)-\mu_{\alpha}(h)$
$=\operatorname{Re}\int\mathrm{e}^{\mathop{\mathrm{i}}x}\,\mathds{P}\left(2^{-1/2}B_{1}\in\mathrm{d}x\right)-\operatorname{Re}\int\mathrm{e}^{\mathop{\mathrm{i}}x}\,\mathds{P}\left(\alpha^{-1/\alpha}L_{1}\in\mathrm{d}x\right)$
$=\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}2^{-1/2}B_{1}}\right]-\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\alpha^{-1/\alpha}L_{1}}\right]$
$=\mathrm{e}^{-1/4}-\mathrm{e}^{-1/(2\alpha)},$
which yields
$\liminf_{\alpha\uparrow 2}\frac{\|\mu_{2}-\mu_{\alpha}\|_{\mathrm{TV}}}{2-\alpha}\geq\lim_{\alpha\uparrow 2}\frac{\mathrm{e}^{-1/4}-\mathrm{e}^{-1/(2\alpha)}}{2-\alpha}=\frac{1}{8}\,\mathrm{e}^{-1/4}.\qed$
Acknowledgement.
C.-S. Deng is supported by the National Natural Science Foundation of China (12371149) and the Alexander von Humboldt Foundation. R.L. Schilling is supported by the 6G-Life project (BMBF Grant No. 16KISK001K) and the ScaDS.AI centre. L. Xu is supported by the National Natural Science Foundation of China No. 12071499, The Science and Technology Development Fund (FDCT) of Macau S.A.R. FDCT 0074/2023/RIA2, and the University of Macau grants MYRG2020-00039-FST, MYRG-GRG2023-00088-FST.

### v2

@@
-Total variation distance between SDEs with stable noise and Brownian motion with applications to Poisson PDEs
+Total variation distance between SDEs with stable noise and Brownian motion
 Abstract.
-We consider a $d$-dimensional stochastic differential equation (SDE) of the form $dU_{t}=b(U_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}Z_{t}$ with initial condition $U_{0}=u\in{\mathds{R}^{d}}$ and additive noise.
-Denote by $X_{t}^{x}$ the (unique) solution with initial value $X_{0}^{x}=x$, if the driving noise $Z_{t}$ is a $d$-dimensional rotationally symmetric $\alpha$-stable ($21<\alpha<$ 2) Lévy process, and by $Y_{t}^{y}$ the (unique) solution with initial value $Y_{0}^{y}=y$, if the driving noise is a $d$-dimensional Brownian motion.
-We give precise estimates of the total variation distance $\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\|_{\mathrm{TV}}$ for all $t>0$, and we show that the ergodic measures $\mu_{\alpha}$ and $\mu_{2}$ of $X_{t}^{x}$ and $Y_{t}^{y}$, respectively, satisfy $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right]$.
-As an application, we consider the following $d$-dimensional Poisson PDE
-$-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
-where $h\in C_{b}({\mathds{R}^{d}},\real)$ and $1<\alpha\leq 2$.
-We show that this PDE admits a solution $f_{\alpha}$, which is continuous, but may have linear growth.
-For fixed $\alpha_{0}\in(1,2)$, we show that for any $\alpha_{0}\leq\alpha<2$ there is a constant $C(\alpha_{0})$ such that
-$\left|f_{\alpha}(x)-f_{2}(x)\right|\leq C(\alpha_{0})(1+|x|)\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha}\quad\text{for all $x$}.$
+We consider a $d$-dimensional stochastic differential equation (SDE) of the form $\mathrm{d}U_{t}=b(U_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}Z_{t}$, let $X_{t}$ be the solution if the driving noise $Z_{t}$ is a $d$-dimensional rotationally symmetric $\alpha$-stable process ($1<\alpha<2$), and let $Y_{t}$ be the solution if the driving noise is a $d$-dimensional Brownian motion.
+Continuing the work in [5], we derive an estimate of the total variation distance $\|\mathrm{law}(X_{t})-\mathrm{law}(Y_{t})\|_{\mathrm{TV}}$ for all $t>0$, and we show that the ergodic measures $\mu_{\alpha}$ and $\mu_{2}$ of $X_{t}$ and $Y_{t}$, respectively, satisfy
+$\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
+We shall show that this bound is optimal with respect to $\alpha$ by an Ornstein–Uhlenbeck SDE.
+Combining this bound with a recent interpolation result from [8], we can derive a bound in Wasserstein-$p$ distance ($0<p<1$):
+$\|\mu_{\alpha}-\mu_{2}\|_{W_{p}}\leq\frac{Cd^{(p+3)/2}\log(1+d)}{\alpha-1}(2-\alpha).$
+Key Words: Total variation distance, Wasserstein-$p$ distance, stochastic differential equation, Poisson equation, stable process.
 1. Introduction
-This paper can be seen as a continuation of our previous paper [5], where we studied the 1-Wasserstein distance between stochastic differential equations (SDEs) with different drivers based on the methods developed in [1].
-This means that we shall work in the same set-up and under similar assumptions, which we will briefly recall.
-1.1. General setup and notation
-Consider the SDEs
+Let us consider the SDEs
 (1.1) $\mathrm{d}X_{t}=b(X_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}L_{t},\quad X_{0}=x,$
 (1.2) $\mathrm{d}Y_{t}=b(Y_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}B_{t},\quad Y_{0}=y,$
-driven by a standard $d$-dimensioal Brownian motion $B_{t}$ and a stable Lévy process $L_{t}$ with characteristic function $\mathds{E}\,\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, $21<\alpha<$ 2; this is a standard rotationally symmetric $\alpha$-stable process run at half speed.
-The drift coefficient $b:{\mathds{R}^{d}}\to{\mathds{R}^{d}}$ will be specified further down, $\sigma$ is an invertible $d\times d$ matrix.
+driven by a standard $d$-dimensioal Brownian motion $B_{t}$ and a stable Lévy process $L_{t}$ with characteristic function $\mathds{E}\,\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, $1<\alpha<2$; this is a standard rotationally symmetric $\alpha$-stable process run at half speed.
 Our main assumptions are as follows:
 (H1).
-There exist constants $\theta_{0}>0$ and $K\geq 0$ such that
-$\langle x-y,b(x)-b(y)\rangle\leq-\theta_{0}|x-y|^{2}+K\quad\text{for all\ \ }x,y\in{\mathds{R}^{d}};$
+$\sigma\in\real^{d\times d}$ is an invertible $d\times d$ matrix.
 (H2).
-$b\in C^{2}({\mathds{R}^{d}},{\mathds{R}^{d}})$ and there exist constants $\theta_{1},\theta_{2}\geq 0$ such that for all $x,v,w\in{\mathds{R}^{d}}$,
+$b\in C^{2}({\mathds{R}^{d}},{\mathds{R}^{d}})$ and there exist constants $\theta_{0}>0$, $K\geq 0$, $\theta_{1}\geq 0$, $\theta_{2}\geq 0$ such that for all $x,y,v,v_{1},v_{2}\in{\mathds{R}^{d}}$,
+(1.3) $\langle x-y,b(x)-b(y)\rangle$ $\leq-\theta_{0}|x-y|^{2}+K,$
 (1.4) $|\nabla_{v}b(x)|$ $\leq\theta_{1}|v|,$
-(1.4) $|\nabla_{w}\nabla_{v}b(x)|$ $\leq\theta_{2}|v||w|.$
-If $f:{\mathds{R}^{d}}\to\real$ is smooth enough, the directional derivative in direction $v\in{\mathds{R}^{d}}$ is defined as
+(1.5) $|\nabla_{v_{2}}\nabla_{v_{1}}b(x)|$ $\leq\theta_{2}|v_{1}||v_{2}|.$
+It is well known that under (1.4) the both SDEs (1.1) and (1.2) have unique non-explosive (strong) solutions, which we denote by $X^{x}_{t}$ and $Y^{y}_{t}$ respectively.
+From the classical Lyapunov function criterion, see e.g.
+[11], we know that under (H1) and (H2) the solutions to the SDEs (1.1) and (1.2) are ergodic.
+Denote by $\mu_{\alpha}$ and $\mu_{2}$ the respective ergodic measures of $X^{x}_{t}$ and $Y^{y}_{t}$.
+Continuing the work in Continuing the work in [5], we establish an optimal bound between $X_{t}^{x}$ and $Y_{t}^{y}$ in total variation distance and consequently an optimal bound between $\mu_{\alpha}$ and $\mu_{2}$ as $t\rightarrow\infty$.
+Combining this bound with a recent interpolation result from [8], we can extend these bounds to the Wasserstein-$p$ distance with $0<p<1$.
+1.1. Notation
+Before giving our main results, let us first introduce the notation that will be used in sequel.
+We denote by $C({\mathds{R}^{d}},\real^{m})$ the set of all continuous functions from $\mathds{R}^{d}$} to m and by $C_{b}({\mathds{R}^{d}},\real)$ the set of all bounded continuous functions from $\mathds{R}^{d}$} to .
+Given $k\in\mathds{N}$, denote by $C^{k}({\mathds{R}^{d}},\real)$ the set of all continuous functions from $\mathds{R}^{d}$} to with continuous 11st,…,$k$-th order derivatives, and by $C^{k}_{b}({\mathds{R}^{d}},\real)$ the set of all bounded continuous functions from $\mathds{R}^{d}$} to with bounded continuous 11st,…,$k$-th order derivatives.
+Let $f:{\mathds{R}^{d}}\to\real$ be sufficiently smooth, the directional derivative in direction $v\in{\mathds{R}^{d}}$ is defined as
 $\nabla_{v}f(x):=\lim_{\epsilon\to 0}\frac{f(x+\epsilon v)-f(x)}{\epsilon}=\nabla f(x)\cdot v$
-and $\nabla_{w}\nabla_{v}f(x):=\nabla_{w}(\nabla_{v}f)(x)=\nabla^{2}f(x)v\cdot w$ for $v,w\in{\mathds{R}^{d}}$. The vector and matrix norms of $\nabla f$ and $\nabla^{2}f$ are given by
-$|\nabla f(x)|=\sup_{v\in{\mathds{R}^{d}},|v|=1}\left|\nabla_{v}f(x)\right|\quad\text{and}\quad\|\nabla f(x)\|_{\mathrm{op}}=\sup_{v,w\in{\mathds{R}^{d}},|v|=|w|=1}\left|\nabla_{w}\nabla_{v}f(x)\right|$
-[…]
+and $\nabla_{v_{2}}\nabla_{v_{1}}f(x):=\nabla_{v_{2}}(\nabla_{v_{1}}f)(x)=\nabla^{2}f(x){v_{1}}\cdot{v_{2}}$ for $v_{1},v_{2}\in{\mathds{R}^{d}}$.
+The vector and matrix norms of $\nabla f$ and $\nabla^{2}f$ are given by
+$|\nabla f(x)|=\sup_{v\in{\mathds{R}^{d}},|v|=1}\left|\nabla_{v}f(x)\right|$
+and
+$\|\nabla^{2}f(x)\|_{\mathrm{op}}=\sup_{v_{1},v_{2}\in{\mathds{R}^{d}},|v_{1}|=|v_{2}|=1}\left|\nabla_{v_{2}}\nabla_{v_{1}}f(x)\right|=\sup_{v_{1},v_{2}\in{\mathds{R}^{d}},|v_{1}|=|v_{2}|=1}\left\langle\nabla^{2}f(x),v_{1}v_{2}^{\top}\right\rangle_{\rm HS}.$
+Moreover, we define
+$\|\nabla f\|_{\infty}:=\sup_{x\in{\mathds{R}^{d}}}\|\nabla f(x)\|_{\mathrm{op}},\ \ \ \|\nabla^{2}f\|_{\mathrm{op},\infty}:=\sup_{x\in{\mathds{R}^{d}}}\|\nabla^{2}f(x)\|_{\mathrm{op}}.$
+This notation extends naturally to vector-valued functions $f=(f_{1},f_{2},\cdots,f_{d})^{\top}:{\mathds{R}^{d}}\to{\mathds{R}^{d}}$:
 $\nabla f(x)v$ $=\nabla_{v}f(x)=(\nabla_{v}f_{1}(x),\dots,\nabla_{v}f_{d}(x))^{\top},$
-$\nabla^{2}f(x)vu$ $=\nabla_{u}\nabla_{v}f(x)=(\nabla_{u}\nabla_{v}f_{1}(x),\dots,\nabla_{u}\nabla_{v}f_{d}(x))^{\top}.$
-The Hilbert–Schmidt inner product of two matrices $A,B\in{}^{d\times d}$ is $\left\langle A,B\right\rangle_{\mathrm{HS}}=\sum_{i,j=1}^{d}A_{ij}B_{ij}$, and the Hilbert-Schmidt norm is $\|A\|_{\mathrm{HS}}=\sqrt{\sum_{i,j=1}^{d}A_{ij}^{2}}$.
-Finally, $\mathrm{Lip}(\lambda)$ is the family of Lipschitz functions $f:{\mathds{R}^{d}}\to\real$ with Lipschitz constant $\lambda$ and norm $\|f\|_{\mathrm{Lip}}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|}\leq\lambda$.
-We will frequently need the constants
-(1.5) $A(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{d/2}\Gamma(1-\frac{\alpha}{2})},\quad\omega_{d-1}:=\frac{2\pi^{d/2}}{\Gamma\left(\frac{d}{2}\right)},$
-which can be estimated as follows (see [5, Lemma 3.3]):
-(1.6) $\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}\leq C\quad\text{and}\quad\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\leq C(2-\alpha)\log(1+d).$
-Under (1.3) the both SDEs (1.1) and (1.2) have unique non-explosive (strong) solutions, which we denote by $X_{t}$ and $Y_{t}$, respectively.
-These are time homogeneous Markov processes and write $X_{t}^{x}$ and $Y_{t}^{y}$ to indicate the initial position $X_{0}^{x}=x$ and $Y_{0}^{y}=y$.
-The operator semigroup induced by the Markov process $(X_{t}^{x})_{t\geq 0}$ is
+$\nabla^{2}f(x)v_{1}v_{2}$ $=\nabla_{v_{2}}\nabla_{v_{1}}f(x)=(\nabla_{v_{2}}\nabla_{v_{1}}f_{1}(x),\dots,\nabla_{v_{2}}\nabla_{v_{1}}f_{d}(x))^{\top}.$
+For $x\in{\mathds{R}^{d}}$, we see $\nabla f(x)\in{\mathds{R}^{d}}$ and $\nabla^{2}f(x)\in\real^{d\times d}$.
+The Hilbert–Schmidt inner product of two matrices $A,B\in\real^{d\times d}$ is $\left\langle A,B\right\rangle_{\mathrm{HS}}=\sum_{i,j=1}^{d}A_{ij}B_{ij}$, and the Hilbert-Schmidt norm is $\|A\|_{\mathrm{HS}}=\sqrt{\sum_{i,j=1}^{d}A_{ij}^{2}}$.
+Recall that $X_{t}^{x}$ and $Y_{t}^{y}$ are the solutions to SDEs (1.1) and (1.2) respectively. The operator semigroup induced by the Markov process $(X_{t}^{x})_{t\geq 0}$ is
 $P_{t}f(x)=\mathds{E}f(X^{x}_{t}),\quad f\in C_{b}({\mathds{R}^{d}},\real),\,t>0.$
@@
 In the same way we can consider the semigroup $Q_{t}$ of $(Y_{t}^{y})_{t\geq 0}$ and its generator $\left(\mathscr{A}^{Q},\mathcal{D}(\mathscr{A}^{Q})\right)$.
-From the classical Lyapunov function criterion, see e.g.
-[10], we know that under (H1) and (H2) the solutions to the SDEs (1.1) and (1.2) are ergodic.
-Denote by $\mu_{\alpha}$ and $\mu_{2}$ the respective ergodic measures of $X_{t}^{x}$ and $Y_{t}^{y}$.
-1.2. Main results
-We will now describe the principal results of our paper.
-Write $\mathscr{C}_{b}(1)=\{h\in C_{b}({\mathds{R}^{d}},\real)\>;\>\|h\|_{\infty}\leq 1\}\}$ for the unit ball in $C_{b}({\mathds{R}^{d}},\real)$.
+Write $\mathscr{C}_{b}(1)=\{h\in C_{b}({\mathds{R}^{d}},\real)\>;\>\|h\|_{\infty}\leq 1\}$ for the unit ball in $C_{b}({\mathds{R}^{d}},\real)$.
 The total variation distance of two probability measures (on $\mathds{R}^{d}$}) $\mu$ and $\nu$ is given by
 $\|\mu-\nu\|_{\mathrm{TV}}:=\sup_{h\in\mathscr{C}_{b}(1)}|\mu(h)-\nu(h)|$
-(as usual we write $\mu(h):=\int h\,\mathrm{d}\mu$ etc.). It is clear that $\|\mu-\nu\|_{\mathrm{TV}}\leq 2$ for all probability measures $\mu$ and $\nu$.
+(as usual we write $\mu(h):=\int h\,\mathrm{d}\mu$ etc.).
+It is clear that $\|\mu-\nu\|_{\mathrm{TV}}\leq 2$ for all probability measures $\mu$ and $\nu$.
+For $p\in(0,1]$, the Wasserstein-$p$ distance for two probability measures $\mu,\nu$ is given by
+$\|\mu-\nu\|_{W_{p}}=\sup_{|h(x)-h(y)|\leq|x-y|^{p}}|\mu(h)-\nu(h)|.$
+Finally, $\mathrm{Lip}(1)$ is the family of Lipschitz functions $f:{\mathds{R}^{d}}\to\real$ with Lipschitz constant $1$ and norm $\|f\|_{\mathrm{Lip}}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|}\leq 1$.
+1.2. Main results
 Throughout this paper, $C,C_{1},C_{2}$ denote positive constants which may depend on $\theta_{0},\theta_{1},\theta_{2},K,\|\sigma\|_{\mathrm{op}},\|\sigma^{-1}\|_{\mathrm{op}}$, but they are always independent of $d$ and $\alpha$; their values may change from line to line.
+Recall that $\mu_{\alpha}$ and $\mu_{2}$ are respective ergodic measures of $X^{x}_{t}$ and $Y^{y}_{t}$ respectively.
 Theorem 1.1.
@@
 $\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\|_{\mathrm{TV}}$
-$\qquad\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|+\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
+$\qquad\leq C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|+\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
 In particular,
-(1.7) $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right].$
-In the estimate (1.7) we cannot hope, in general, for a faster speed of convergence than $2-\alpha$. This can be seen from the one-dimensional Ornstein–Uhlenbeck, cf. Proposition 7.1 below.
-We can apply this theorem to study the convergence rate of Poisson PDEs as follows ($1<\alpha\leq 2$):
-(1.8) $-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
-where $h\in C_{b}({\mathds{R}^{d}},\real)$,
-$-(-\Delta)^{\alpha/2}f(x)=\int_{{\mathds{R}^{d}}\setminus\{0\}}\left[f(x+z)-f(x)-\langle\nabla f(x),z\rangle\mathds{1}_{(0,1)}(|z|)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
-is the fractional Laplacian for $\alpha\in(1,2)$, and $\Delta f(x)=\sum_{i=1}^{d}\partial^{2}_{i}f(x)$ is the usual Laplace operator.
-This type of Poisson equation has many applications in central limit theorems related to SDEs and approximations of probability measures, see [9] and [7, 20].
-Our second theorem is as follows:
-Theorem 1.2.
-Assume (H1) and (H2).
-  1. a)
-    For any $\alpha\in(1,2]$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the equation (1.8) admits a solution $f_{\alpha}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$.
-  2. b)
-    Let $\alpha_{0}\in(1,2)$ be arbitrary. For any $\alpha\in[\alpha_{0},2)$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the solution $f_{\alpha}$ from a) satisfies
-    $\|f_{\alpha}-f_{2}\|_{\mathrm{lin}}\leq C(\alpha_{0})\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha},$
-    with a constant $C(\alpha_{0})$ depending on $\alpha_{0},\theta_{0},\theta_{1},\theta_{2},K$.
-2. The key gradient estimate for Theorem 1.1
-This section is devoted to proving the following gradient estimate, which is crucial for the proof of our main result.
+(1.6) $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
+We shall show that the bound in (1.6) is optimal with respect to $\alpha$ by a one-dimensional Ornstein–Uhlenbeck system, cf. Proposition 5.1 below.
+Applying a recent interpolation result established in [8], we obtain the following corollary.
+Corollary 1.2.
+Assume (H1) and (H2). For any $p\in(0,1)$ and $\alpha\in(1,2)$,
+$\|\mu_{\alpha}-\mu_{2}\|_{W_{p}}\leq\frac{Cd^{(p+3)/2}\log(1+d)}{\alpha-1}(2-\alpha).$
+2. A key gradient estimate for Theorem 1.1
+This section is devoted to proving the following gradient estimate by an integration by parts in Malliavin calculus, which is crucial for the proof of our main result.
 Proposition 2.1.
-Assume (H2). For any $t\in(0,1]$, $\alpha\in(0,2)$ and $f\in C^{2}_{b}({\mathds{R}^{d}},\real)$,
-[…]
+Assume (1.4) and (1.5). For any $t\in(0,1]$, $\alpha\in(0,2)$ and $f\in C^{2}_{b}({\mathds{R}^{d}},\real)$,
+$\|P_{t}\nabla^{2}f\|_{\mathrm{op},\infty}\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\|\nabla f\|_{\infty}.$
 2.1. A random time-change method
@@
 Let $\gamma^{\epsilon}$ be the inverse function of $\ell^{\epsilon}$, i.e., $\ell^{\epsilon}\circ\gamma^{\epsilon}_{t}=t$ for all $t\geq\ell_{0}^{\epsilon}$ and $\gamma^{\epsilon}\circ\ell^{\epsilon}_{t}=t$ for all $t\geq 0$.
-By construction, $t\mapsto\gamma_{t}^{\epsilon}$ is absolutely continuous on $[\ell_{0}^{\epsilon},\infty)$.
+By construction, $t\mapsto\gamma_{t}^{\epsilon}$ is absolutely continuous on $\ell_{0}^{\epsilon},\infty)$.
 Define
@@
 For fixed $t>\ell^{\epsilon}_{0}$, consider the Hilbert space $H:=L^{2}([\ell^{\epsilon}_{0},t];{\mathds{R}^{d}})$ equipped with the inner product
-$\langle f,g\rangle_{H}:=\int_{\ell^{\epsilon}_{0}}^{t}\langle f(s),g(s)\rangle\,\mathrm{d}s\quad\text{for $f,g\in H$}.$
+[…]
 Let $W=\{W(h)\,;\,h\in H\}$ be an isonormal Gaussian process associated with $H$, i.e.
@@
 The following auxiliary lemma is crucial for proving Proposition 2.1.
-Lemma 2.2.
-Assume (H2), and fix $\ell\in\mathds{S}$ and $\epsilon\in(0,1)$. For any $t>0$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
-$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell^{\epsilon}}_{t}\right)\right]\right|\leq C\left(\frac{1}{\sqrt{\ell^{\epsilon}_{t}-\ell^{\epsilon}_{0}}}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell^{\epsilon}}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
+Lemma $2.$
+Assume (1.4) and (1.5), and fix $\ell\in\mathds{S}$ and $\epsilon\in(0,1)$. For any $v\in\real^{d}$, $t>0$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
+$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell^{\epsilon}}_{t}\right)\right]\right|\leq C|v|\left(\frac{1}{\sqrt{\ell^{\epsilon}_{t}-\ell^{\epsilon}_{0}}}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell^{\epsilon}}_{t})\right|^{2}\right]}.$
 Proof.
@@
 $J_{s,t}:=\exp\left(\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right),\quad\ell^{\epsilon}_{0}\leq s\leq t.$
-For any fixed $t>\ell^{\epsilon}_{0}$ and $1\leq i\leq d$, let
-(2.7) $u_{i}(s):=\frac{1}{t-\ell^{\epsilon}_{0}}\,\sigma^{-1}J_{s,t}^{-1}e_{i}.$
+For any fixed $t>\ell^{\epsilon}_{0}$ and $v\in\real^{d}$, let
+(2.7) $u_{0}(s):=\frac{1}{t-\ell^{\epsilon}_{0}}\,\sigma^{-1}J_{s,t}^{-1}v.$
 It can be easily verified that
-$D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}=e_{i},\quad 1\leq i\leq d.$
-Combining this with the duality (2.5) between the Malliavin derivative and the divergence operator gives us, for any $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $1\leq i\leq d$,
-$\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]$ $=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),e_{i}\big{\rangle}\right]=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}\big{\rangle}\right]$
-$=\mathds{E}^{\mathds{W}}\left[D_{u_{i}}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]=\mathds{E}^{\mathds{W}}\left[f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\delta\left(u_{i}\right)\right].$
-By the Cauchy–Schwarz inequality and [12, Proposition 1.3.1], we obtain
-(2.8) $\begin{split}\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|&\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\delta(u_{i})^{2}\right]}\\ &\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\|u_{i}\|_{H}^{2}+\|Du_{i}\|_{H\otimes H}^{2}\right]},\end{split}$
+$D_{u_{0}}Z^{x;\ell^{\epsilon}}_{t}=v.$
+Combining this with the duality (2.5) between the Malliavin derivative and the divergence operator gives us, for any $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$,
+$\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]$ $=\mathds{E}^{\mathds{W}}\left[\big\langle\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),v\big\rangle\right]=\mathds{E}^{\mathds{W}}\left[\big\langle\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),D_{u_{0}}Z^{x;\ell^{\epsilon}}_{t}\big\rangle\right]$
+$=\mathds{E}^{\mathds{W}}\left[D_{u_{0}}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]=\mathds{E}^{\mathds{W}}\left[f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\delta\left(u_{0}\right)\right].$
+By the Cauchy–Schwarz inequality and [13, Proposition 1.3.1], we obtain
+(2.8) $\begin{split}\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|&\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\delta(u_{0})^{2}\right]}\\ &\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\|u_{0}\|_{H}^{2}+\|Du_{0}\|_{H\otimes H}^{2}\right]},\end{split}$
 where $\|\cdot\|_{H\otimes H}$ denotes the norm on the tensor space $H\otimes H$. By (1.4) and (2.7), one has
-(2.9) $\begin{split}\|u_{i}\|_{H}^{2}&=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\exp\left(-\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)e_{i}\right|^{2}\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\exp\left(2\theta_{1}\int_{s}^{t}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{t-\ell^{\epsilon}_{0}}\,\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}.\end{split}$
-Similarly, using (1.3) and (2.6) gives for any $v\in H$
-(2.10) $\left|D_{v}Z^{x;\ell^{\epsilon}}_{t}\right|\leq\|\sigma\|_{\mathrm{op}}\|v\|_{H}\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{t-\ell^{\epsilon}_{0}}.$
+(2.9) $\begin{split}\|u_{0}\|_{H}^{2}&=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\exp\left(-\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)v\right|^{2}\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}|v|^{2}}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\exp\left(2\theta_{1}\int_{s}^{t}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}|v|^{2}}{t-\ell^{\epsilon}_{0}}\,\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}.\end{split}$
+Similarly, using (1.4) and (2.6) gives for any $u\in H$
+(2.10) $\left|D_{u}Z^{x;\ell^{\epsilon}}_{t}\right|\leq\|\sigma\|_{\mathrm{op}}\|u\|_{H}\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{t-\ell^{\epsilon}_{0}}.$
 Furthermore, by the chain rule for matrix exponential functions (see e.g. [14]), one has
-$\begin{split}D_{v}J_{s,t}^{-1}&=\int_{0}^{1}\exp\left(-\tau\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\left(-\int_{s}^{t}\nabla^{2}b\left(Z^{x;\ell^{\epsilon}}_{r}\right)D_{v}Z^{x;\ell^{\epsilon}}_{r}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\mbox{}\times\exp\left(-(1-\tau)\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}\tau,\end{split}$
+$\begin{split}D_{u}J_{s,t}^{-1}&=\int_{0}^{1}\exp\left(-\tau\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\left(-\int_{s}^{t}\nabla^{2}b\left(Z^{x;\ell^{\epsilon}}_{r}\right)D_{u}Z^{x;\ell^{\epsilon}}_{r}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\mbox{}\times\exp\left(-(1-\tau)\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}\tau,\end{split}$
 which, together with (1.4), (1.5) and (2.10), yields that
-$\|D_{v}J_{s,t}^{-1}\|_{\mathrm{op},\infty}\leq\|\sigma\|_{\mathrm{op}}\theta_{2}\|v\|_{H}\sqrt{t-\ell^{\epsilon}_{0}}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}},\quad s\in[\ell^{\epsilon}_{0},t].$
-Then, it follows from the chain rule and (2.7) that
-$\|D_{v}u_{i}\|^{2}_{H}$ $=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}}^{t}\left|\sigma^{-1}\left\{D_{v}J_{s,t}^{-1}\right\}e_{i}\right|^{2}\mathrm{d}s$
-$\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}\|v\|_{H}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
-Since this holds for any $v\in H$, one can deduce that
-$\|Du_{i}\|^{2}_{H\otimes H}\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
+$\|D_{u}J_{s,t}^{-1}\|_{\mathrm{op},\infty}\leq\|\sigma\|_{\mathrm{op}}\theta_{2}\|u\|_{H}\sqrt{t-\ell^{\epsilon}_{0}}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}},\quad s\in[\ell^{\epsilon}_{0},t].$
+Then, it follows from (2.7) that
+$\|D_{u}u_{0}\|^{2}_{H}$ $=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\left\{D_{u}J_{s,t}^{-1}\right\}v\right|^{2}\mathrm{d}s$
+$\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}\|u\|_{H}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}|v|^{2}.$
+Since this holds for any $u\in H$, one can deduce that
+$\|Du_{0}\|^{2}_{H\otimes H}\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}|v|^{2}.$
 Combining this with (2.8) and (2.9) gives us for $t\geq\ell^{\epsilon}_{0}$,
-$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|$
-$\quad\leq\left(\frac{\|\sigma^{-1}\|_{\mathrm{op}}}{\sqrt{t-\ell^{\epsilon}_{0}}}\,\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}+\|\sigma^{-1}\|_{\mathrm{op}}\|\sigma\|_{\mathrm{op}}\theta_{2}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}$
-$\quad\leq C\left(\frac{1}{\sqrt{t-\ell^{\epsilon}_{0}}}+\gamma^{\epsilon}_{t}\right)\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}.$
+$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|$
+$\quad\leq\left(\frac{\|\sigma^{-1}\|_{\mathrm{op}}|v|}{\sqrt{t-\ell^{\epsilon}_{0}}}\,\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}+\|\sigma^{-1}\|_{\mathrm{op}}\|\sigma\|_{\mathrm{op}}\theta_{2}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}|v|\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big|^{2}\right]}$
+$\quad\leq C|v|\left(\frac{1}{\sqrt{t-\ell^{\epsilon}_{0}}}+\gamma^{\epsilon}_{t}\right)\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big|^{2}\right]}.$
 Since $Z_{\ell_{t}^{\epsilon}}^{x;\ell^{\epsilon}}=X_{t}^{x;\ell^{\epsilon}}$, we can replace $t$ by $\ell_{t}^{\epsilon}$ and get the desired estimate. ∎
@@
 Lemma 2.3.
-Assume (H2). For any $\alpha\in(0,2)$, $t\in(0,1]$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
-[…]
+Assume (1.4) and (1.5). For any $\alpha\in(0,2)$, $t\in(0,1]$ and $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$,
+$\left\|P_{t}\nabla f\right\|_{\infty}\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\|f\|_{\infty}.$
 Proof.
-Letting $\epsilon\downarrow 0$ in Lemma 2.2 and using $\ell^{\epsilon}_{t}\downarrow\ell_{t}$ and [22, Lemma 2.2], we get for $t>0$,
-$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\right|\leq C\left(\ell_{t}^{-1/2}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
-Now we “unfreeze” the fixed path $\ell$ and use (2.2). From the Cauchy–Schwarz inequality we get for $t>0$,
-$\left|P_{t}(\partial_{i}f)(x)\right|$ $=\left|\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)\right|$
-$\leq\int_{\mathds{S}}\big{|}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\big{|}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
-$\leq C\mathrm{e}^{2\theta_{1}t}\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
-$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)^{2}\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)}$
-$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1}+t^{2}\right)\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\mathds{E}\left[\big{|}f(X^{x}_{t})\big{|}^{2}\right]}$
-$=C\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}\left[S_{t}^{-1}\right]+t^{2}}\,\sqrt{P_{t}|f|^{2}(x)}$
-$\leq C\mathrm{e}^{2\theta_{1}t}\left(\sqrt{\mathds{E}\left[S_{t}^{-1}\right]}+t\right)\sqrt{P_{t}|f|^{2}(x)}.$
+Letting $\epsilon\downarrow 0$ in Lemma 2.2 and using $\ell^{\epsilon}_{t}\downarrow\ell_{t}$ and [24, Lemma 2.2], we get for $t>0$,
+$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\right|\leq C|v|\left(\ell_{t}^{-1/2}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell}_{t})\right|^{2}\right]}.$
+Now we “unfreeze” the fixed path $\ell$ and use (2.2). From the Cauchy–Schwarz inequality we get for all $t>0$, $v\in\real^{d}$ and $x\in\real^{d}$,
+$\left|P_{t}(\nabla_{v}f)(x)\right|$ $=\left|\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)\right|$
+$\leq\int_{\mathds{S}}\big|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\big|\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
+$\leq C|v|\mathrm{e}^{2\theta_{1}t}\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f(X^{x;\ell}_{t})\big|^{2}\right]}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
+$\leq C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)^{2}\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\big|f(X^{x;\ell}_{t})\big|^{2}\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)}$
+$\leq C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1}+t^{2}\right)\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\mathds{E}\left[\big|f(X^{x}_{t})\big|^{2}\right]}$
+$=C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}\left[S_{t}^{-1}\right]+t^{2}}\,\sqrt{P_{t}|f|^{2}(x)}$
+$\leq C|v|\mathrm{e}^{2\theta_{1}t}\left(\sqrt{\mathds{E}\left[S_{t}^{-1}\right]}+t\right)\|f\|_{\infty}.$
 From [4, Lemma 4.1] we know the following moment identity for $\alpha\in(0,2)$
@@
 Proof of Proposition 2.1.
-It follows from Lemma 2.3 that for all $t\in(0,1]$, $\alpha\in(0,2)$ and $x\in{\mathds{R}^{d}}$,
-$\left\|\mathds{E}\left[\nabla^{2}f(X^{x}_{t})\right]\right\|_{\mathrm{HS}}^{2}$ $=\sum_{i,j=1}^{d}\left|\mathds{E}\left[\partial_{i}\partial_{j}f(X^{x}_{t})\right]\right|^{2}$
-[…]
-[…]
+It follows from Lemma 2.3 that for all $t\in(0,1]$ and $\alpha\in(0,2)$,
+$\|P_{t}\nabla^{2}f\|_{\mathrm{op},\infty}$ $=\sup_{x\in\real^{d}}\sup_{v_{1},v_{2}\in\real^{d},|v_{1}|=|v_{2}|=1}\left|P_{t}\nabla_{v_{2}}\nabla_{v_{1}}f(x)\right|$
+$\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\sup_{v_{1}\in\real^{d},|v_{1}|=1}\|\nabla_{v_{1}}f\|_{\infty}.$
 Finally,
-$\sum_{j=1}^{d}P_{t}|\partial_{j}f|^{2}(x)=P_{t}|\nabla f|^{2}(x)\leq\|\nabla f\|_{\infty}^{2},$
+$\sup_{v\in\real^{d},|v|=1}\|\nabla_{v}f\|_{\infty}=\sup_{x\in\real^{d}}\sup_{v\in\real^{d},|v|=1}\left|\langle\nabla f(x),v\rangle\right|=\|\nabla f\|_{\infty},$
 and this completes the proof. ∎
@@
 This completes the proof of (3.2). ∎
+In the following lemma, we will frequently need the constants
+$A(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{d/2}\Gamma(1-\frac{\alpha}{2})},\quad\omega_{d-1}:=\frac{2\pi^{d/2}}{\Gamma\left(\frac{d}{2}\right)},$
+which can be estimated as follows (see [5, Lemma 3.3]):
+(3.3) $\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}\leq C\quad\text{and}\quad\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\leq C(2-\alpha)\log(1+d).$
 Lemma 3.2.
@@
     $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}$
-    $\qquad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d.$
+    $\qquad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d\log(1+d)+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d.$
   2. b)
     For any $\alpha\in(1,2)$, $0<s<t\leq 1$ and $k=1,2,\dots$,
-    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\leq C_{1}d^{3/2}(2-\alpha)\left((t-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{e}^{-C_{2}k}.$
+    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\leq C_{1}(2-\alpha)\left((t-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{e}^{-C_{2}k}d\log(1+d).$
 Proof.
@@
 $\left(\mathscr{A}^{P}-\mathscr{A}^{Q}\right)f(x)$
-=$=\left\{\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
+$=\left\{\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
 $\quad\mbox{}+\int_{|z|\geq 1}\left[f(x+\sigma z)-f(x)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
@@
 $=\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x+r\sigma z),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}$
-=$=\left\{\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
+$=\left\{\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
 $\qquad\mbox{}+\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x+r\sigma z)-\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
@@
 Combining this with (3.3), Proposition 2.1 and the fact that $\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\leq\sqrt{d}\|\sigma\sigma^{\top}\|_{\mathrm{op}}\leq\sqrt{d}\|\sigma\|_{\mathrm{op}}^{2}$, we obtain for all $\alpha\in(1,2)$,
-(3.4) $\begin{split}|P_{t-s}\mathsf{J}_{11}(x)|&=\left|\frac{1}{2}\left[\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right]\big{\langle}P_{t-s}\nabla^{2}f(x),\sigma\sigma^{\top}\big{\rangle}_{\mathrm{HS}}\right|\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{HS}}\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\\ &\leq C(2-\alpha)(t-s)^{-1/\alpha}\|\nabla f\|_{\infty}d\log(1+d).\end{split}$
-Let us now estimate $P_{t-s}\mathsf{J}_{12}(x)$. Since $\|(\sigma z)(\sigma z)^{\top}\|_{\mathrm{HS}}=|\sigma z|^{2}$, it follows from Proposition 2.1 that for any $r\in[0,1]$ and $\alpha\in(1,2)$,
+(3.5) $\begin{split}|P_{t-s}\mathsf{J}_{11}(x)|&=\left|\frac{1}{2}\left[\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right]\big\langle P_{t-s}\nabla^{2}f(x),\sigma\sigma^{\top}\big\rangle_{\mathrm{HS}}\right|\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{HS}}\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|d\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{op},\infty}\|\sigma\|_{\mathrm{op}}^{2}\\ &\leq C(2-\alpha)(t-s)^{-1/\alpha}\|\nabla f\|_{\infty}d\log(1+d).\end{split}$
+Let us now estimate $P_{t-s}\mathsf{J}_{12}(x)$. It follows from Proposition 2.1 that for any $r\in[0,1]$ and $\alpha\in(1,2)$,
 $\left|\left\langle P_{t-s}\left[\nabla^{2}\left(f(x+r\sigma z)-f(x)\right)\right],(\sigma z)(\sigma z)^{\top}\right\rangle_{\mathrm{HS}}\right|$
-$\quad\leq\left\|P_{t-s}\nabla^{2}\{f(\cdot+r\sigma z)-f(\cdot)\}(x)\right\|_{\mathrm{HS}}\left\|(\sigma z)(\sigma z)^{\top}\right\|_{\mathrm{HS}}$
-$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\|\nabla\{f(\cdot+r\sigma z)-f(\cdot)\}\right\|_{\infty}|\sigma z|^{2}$
-$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{[2\|\nabla f\|_{\infty}]\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}r|\sigma z|]\right\}|\sigma z|^{2}$
-$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\right\}|z|^{2}.$
+$\quad\leq\left\|P_{t-s}\nabla^{2}\{f(\cdot+r\sigma z)-f(\cdot)\}(x)\right\|_{\mathrm{op}}|\sigma z|^{2}$
+$\quad\leq C(t-s)^{-1/\alpha}\left\|\nabla\{f(\cdot+r\sigma z)-f(\cdot)\}\right\|_{\infty}|\sigma z|^{2}$
+$\quad\leq C(t-s)^{-1/\alpha}\left\{[2\|\nabla f\|_{\infty}]\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}r|\sigma z|]\right\}|\sigma z|^{2}$
+$\quad\leq C(t-s)^{-1/\alpha}\left\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\right\}|z|^{2}.$
 This yields for $\alpha\in(1,2)$,
-(3.5) $\begin{split}&|P_{t-s}\mathsf{J}_{12}(x)|\\ &\leq\int_{|z|<1}\int_{0}^{1}\left|\langle P_{t-s}\nabla^{2}f(x+r\sigma z)-P_{t-s}\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}\right|(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left(\int_{0}^{1}(1-r)\,\mathrm{d}r\right)\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\frac{C}{2}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z.\end{split}$
+(3.6) $\begin{split}&|P_{t-s}\mathsf{J}_{12}(x)|\\ &\leq\int_{|z|<1}\int_{0}^{1}\left|\langle P_{t-s}\nabla^{2}f(x+r\sigma z)-P_{t-s}\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}\right|(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq C(t-s)^{-1/\alpha}\int_{|z|<1}\left(\int_{0}^{1}(1-r)\,\mathrm{d}r\right)\big\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big\}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\frac{C}{2}(t-s)^{-1/\alpha}\int_{|z|<1}\big\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z.\end{split}$
 Now we are ready to prove Lemma 3.2.
@@
 Therefore, we obtain from (3.6) that for $\alpha\in(1,2)$,
-$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left\{s^{-1/2}\wedge[s^{-1}|z|]\right\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
-$=C(t-s)^{-1/\alpha}\sqrt{d}\left(\int_{|z|\leq\sqrt{s}}s^{-1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z+\int_{\sqrt{s}<|z|<1}s^{-1/2}\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z\right)$
-$=C(t-s)^{-1/\alpha}\,s^{-1/2}\sqrt{d}\,A(d,\alpha)\omega_{d-1}\left(s^{-1/2}\int_{0}^{\sqrt{s}}r^{2-\alpha}\,\mathrm{d}r+\int_{\sqrt{s}}^{1}r^{1-\alpha}\,\mathrm{d}r\right)$
-$=C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\sqrt{d}\,\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
-$\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}.$
+$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C(t-s)^{-1/\alpha}\int_{|z|<1}\left\{s^{-1/2}\wedge[s^{-1}|z|]\right\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
+$=C(t-s)^{-1/\alpha}\left(\int_{|z|\leq\sqrt{s}}s^{-1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z+\int_{\sqrt{s}<|z|<1}s^{-1/2}\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z\right)$
+$=C(t-s)^{-1/\alpha}\,s^{-1/2}A(d,\alpha)\omega_{d-1}\left(s^{-1/2}\int_{0}^{\sqrt{s}}r^{2-\alpha}\,\mathrm{d}r+\int_{\sqrt{s}}^{1}r^{1-\alpha}\,\mathrm{d}r\right)$
+$=C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
+$\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d.$
 Combining all estimates, we get for all $\alpha\in(1,2)$, $x\in{\mathds{R}^{d}}$ and $0<s<t\leq 1$,
@@
 $\quad\leq|P_{t-s}\mathsf{J}_{11}(x)|+|P_{t-s}\mathsf{J}_{12}(x)|+|P_{t-s}\mathsf{J}_{2}(x)|$
-$\quad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left[(2-\alpha)+\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\right]d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
+$\quad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left[(2-\alpha)+\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\right]d\log(1+d)+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
 from which we immediately get the first part of the lemma.
@@
 Using (3.6), we get for all $\alpha\in(1,2)$,
-$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C_{1}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|\leq 1}\mathrm{e}^{-C_{2}k}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
-$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}A(d,\alpha)\omega_{d-1}\int_{0}^{1}r^{2-\alpha}\,\mathrm{d}r$
-$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}$
-$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}.$
+$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C_{1}(t-s)^{-1/\alpha}\int_{|z|\leq 1}\mathrm{e}^{-C_{2}k}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
+$=C_{1}(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}A(d,\alpha)\omega_{d-1}\int_{0}^{1}r^{2-\alpha}\,\mathrm{d}r$
+$=C_{1}(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}$
+$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d.$
 Finally, combining all estimates, we conclude that for all $\alpha\in(1,2)$,
 $\left|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})f(x)\right|$ $\leq|P_{t-s}\mathsf{J}_{11}(x)|+|P_{t-s}\mathsf{J}_{12}(x)|+|P_{t-s}\mathsf{J}_{2}(x)|$
-$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}+\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
+$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d\log(1+d)+\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
 which completes the proof of the second part of the lemma. ∎
-4. Proof of Theorem 1.1
+4. Proofs of Theorem 1.1 of Corollary 1.2
 Lemma 4.1.
@@
 Assume (H1) and (H2). For all $\alpha\in(1,2)$ and $t>0$,
-$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
+$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
 Proof.
@@
 we get
-$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\sup_{x\in{}^{d}}\left|\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h(x)\,\mathrm{d}s\right|.$
+$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\sup_{x\in\real^{d}}\left|\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h(x)\,\mathrm{d}s\right|.$
 From Lemmas 3.2 and 4.1 we see that for all $\alpha\in(1,2)$ and $t\in(0,1]$,
-(4.3) $\begin{split}\sup_{h\in\mathscr{C}_{b}(1)}&\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\leq Cd^{3/2}\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\quad\mbox{}+\frac{Cd}{\alpha-1}(2-\alpha)\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\\ &\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right)+\frac{2Cd}{\alpha-1}(2-\alpha)t^{1/2}\\ &\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).\end{split}$
+(4.3) $\begin{split}\sup_{h\in\mathscr{C}_{b}(1)}&\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\leq Cd\log(1+d)\cdot\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\quad\mbox{}+\frac{Cd}{\alpha-1}(2-\alpha)\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\\ &\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right)+\frac{2Cd}{\alpha-1}(2-\alpha)t^{1/2}\\ &\leq\frac{C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).\end{split}$
 Combining this with
-$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s,$
+$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s,$
 proves assertion for $t\in(0,1]$.
@@
 we obtain
-(4.4) $\begin{split}\sup_{x\in{}^{d}}&\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\\ &\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sum_{k=1}^{\lfloor t\rfloor-1}\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\right\|_{\infty}\mathrm{d}s,\end{split}$
+(4.4) $\begin{split}\sup_{x\in\real^{d}}&\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\\ &\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sum_{k=1}^{\lfloor t\rfloor-1}\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\right\|_{\infty}\mathrm{d}s,\end{split}$
 where we use the ‘empty-sum-convention’ $\sum_{k=1}^{0}\dots=0$. Using (4.3) with $t=1$, we see that for all $\alpha\in(1,2)$,
-$\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
+$\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
 By Lemma 3.2.b), we find for all $\alpha\in(1,2)$,
 $\sup_{h\in\mathscr{C}_{b}(1)}$ $\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s$
-$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\int_{0}^{1}\left((1-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
-$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
-$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k},$
+$\leq C_{1}(2-\alpha)\mathrm{e}^{-C_{2}k}\int_{0}^{1}\left((1-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s\cdot d\log(1+d)$
+$=C_{1}(2-\alpha)\mathrm{e}^{-C_{2}k}\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)d\log(1+d)$
+$\leq\frac{3C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k},$
 and
 $\sup_{h\in\mathscr{C}_{b}(1)}$ $\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+\lfloor t\rfloor}h\right\|_{\infty}\mathrm{d}s$
-$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\int_{0}^{t-\lfloor t\rfloor}\left((t-\lfloor t\rfloor-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
-$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\left((t-\lfloor t\rfloor)^{1-1/\alpha}\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}(t-\lfloor t\rfloor)\right)$
-$\leq C_{1}d^{3/2}(2-\alpha)\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
-$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha).$
+$\leq C_{1}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\int_{0}^{t-\lfloor t\rfloor}\left((t-\lfloor t\rfloor-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s\cdot d\log(1+d)$
+$=C_{1}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\left((t-\lfloor t\rfloor)^{1-1/\alpha}\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}(t-\lfloor t\rfloor)\right)d\log(1+d)$
+$\leq C_{1}(2-\alpha)\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)d\log(1+d)$
+$\leq\frac{3C_{1}d\log(1+d)}{\alpha-1}(2-\alpha).$
 Substituting these estimates into (4.4), shows that for all $t\geq 1$ and $\alpha\in(1,2)$,
-$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}$ $\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\left(1+\sum_{k=1}^{\lfloor t\rfloor-1}\mathrm{e}^{-C_{2}k}\right)$
-$\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
+$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}$ $\leq\frac{C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)\left(1+\sum_{k=1}^{\lfloor t\rfloor-1}\mathrm{e}^{-C_{2}k}\right)$
+$\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
 This completes the proof. ∎
@@
 Assume (H1) and (H2). For all $x,y\in{\mathds{R}^{d}}$ and $t>0$,
-$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|\right].$
+$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}\leq C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|.$
 Proof.
@@
 we can let $t\to\infty$ to get (1.6) from the first part. ∎
-5. Moment estimates for Theorem 1.2
-The following two moment estimates for the solutions to the SDEs (1.2) and (1.1) are needed to obtain moment estimates for the ergodic measures.
-Note that under our assumptions on the drift coefficient, the solution $Y_{t}^{y}$ to the SDE (1.2) has finite second moments, but we will need a precise estimate of the second moment.
-Lemma 5.1.
-Assume (H1). The following moment estimate holds for the solution $(Y_{t}^{y})_{t\geq 0}$ to the SDE (1.2):
-$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq\mathrm{e}^{-\theta_{0}t}|y|^{2}+Cd\left(1-\mathrm{e}^{-\theta_{0}t}\right),\quad\text{for all\ \ }y\in{\mathds{R}^{d}},\;t\geq 0.$
-Proof.
-From (H1) we get for any $x\in{\mathds{R}^{d}}$,
-(5.1) $\begin{split}\langle b(x),x\rangle&\leq-\theta_{0}|x|^{2}+\langle b(0),x\rangle+K\\ &\leq-\theta_{0}|x|^{2}+\sqrt{\theta_{0}}|x|\frac{|b(0)|}{\sqrt{\theta_{0}}}+K\\ &\leq-\theta_{0}|x|^{2}+\frac{\theta_{0}}{2}|x|^{2}+\frac{|b(0)|^{2}}{2\theta_{0}}+K\\ &=-\frac{\theta_{0}}{2}|x|^{2}+C_{1},\end{split}$
-where $C_{1}:=|b(0)|^{2}/(2\theta_{0})+K$. By Itô’s formula,
-$|Y_{t}^{y}|^{2}$ $=|y|^{2}+\int_{0}^{t}\left(2\langle Y_{s}^{y},b(Y_{s}^{y})\rangle+\|\sigma\sigma^{\top}\|_{\mathrm{HS}}^{2}\right)\mathrm{d}s+2\int_{0}^{t}Y_{s}^{y}\,\mathrm{d}B_{s}$
-$\leq|y|^{2}+\int_{0}^{t}\left(-\theta_{0}|Y_{s}^{y}|^{2}+2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d\right)\mathrm{d}s+2\int_{0}^{t}Y_{s}^{y}\,\mathrm{d}B_{s}.$
-Using the finiteness of the second moment we get that
-$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]=|y|^{2}+\int_{0}^{t}\mathds{E}\left[2\langle Y_{s}^{y},b(Y_{s}^{y})\rangle+\|\sigma\sigma^{\top}\|_{\mathrm{HS}}^{2}\right]\mathrm{d}s;$
-in particular $\mathds{E}\left[|Y_{t}^{y}|^{2}\right]$ is differentiable in $t$. Thus, the above inequality shows
-$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq-\theta_{0}\mathds{E}\left[|Y_{t}^{y}|^{2}\right]+\left(2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d\right).$
-By the Gronwall inequality, we see that for all $t>0$,
-$\mathds{E}\left[|Y_{t}^{y}|^{2}\right]\leq\mathrm{e}^{-\theta_{0}t}|y|^{2}+\frac{2C_{1}+\|\sigma\|_{\mathrm{op}}^{4}d}{\theta_{0}}\left(1-\mathrm{e}^{-\theta_{0}t}\right).$
-This gives the desired estimate. ∎
-Lemma 5.2.
-Assume (H1). The following moment estimate holds for the solution $(X_{t}^{x})_{t\geq 0}$ to the SDE (1.1) with $\alpha\in(1,2)$:
-$\mathds{E}\left[\left(1+|X_{t}^{x}|^{2}\right)^{1/2}\right]\leq\mathrm{e}^{-\theta_{0}t/2}\left(1+|x|^{2}\right)^{1/2}+\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right)\quad\text{for all\ \ }x\in{\mathds{R}^{d}},\;t\geq 0.$
-Proof.
-Set $f(\cdot):=\sqrt{1+|\cdot|^{2}}$. By Dynkin’s formula,
-$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}f(X_{t}^{x})=\mathds{E}\mathscr{A}^{P}f(X_{t}^{x})=\mathds{E}\left[\mathsf{J}_{1}+\mathsf{J}_{2}+\mathsf{J}_{3}\right],$
-where
-$\mathsf{J}_{1}$ $:=\langle\nabla f(X_{t}^{x}),b(X_{t}^{x})\rangle,$
-$\mathsf{J}_{2}$ $:=\int_{|z|\geq 1}\left[f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z,$
-$\mathsf{J}_{3}$ $:=\int_{|z|<1}\left[f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})-\langle\nabla f(X_{t}^{x}),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z.$
-We will estimate the terms $\mathsf{J}_{i}$, $i=1,2,3$, separately. From (5.1) we get
-$\mathsf{J}_{1}$ $=\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}\langle X_{t}^{x},b(X_{t}^{x})\rangle$
-$\leq\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}\left(-\frac{\theta_{0}}{2}|X_{t}^{x}|^{2}+C_{1}\right)$
-$=-\frac{\theta_{0}}{2}\left(1+|X_{t}^{x}|^{2}\right)^{1/2}+\left(C_{1}+\frac{\theta_{0}}{2}\right)\left(1+|X_{t}^{x}|^{2}\right)^{-1/2}$
-$\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+C_{2},$
-where $C_{2}:=C_{1}+\theta_{0}/2$. Since
-$\langle 2X_{t}^{x}+\sigma z,\sigma z\rangle\leq(|X_{t}^{x}+\sigma z|+|X_{t}^{x}|)|\sigma z|\leq\left(\sqrt{1+|X_{t}^{x}+\sigma z|^{2}}+\sqrt{1+|X_{t}^{x}|^{2}}\right)|\sigma z|,$
-we have
-$f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})=\frac{\langle 2X_{t}^{x}+\sigma z,\sigma z\rangle}{\sqrt{1+|X_{t}^{x}+\sigma z|^{2}}+\sqrt{1+|X_{t}^{x}|^{2}}}\leq|\sigma z|.$
-Therefore,
-$\mathsf{J}_{2}\leq\|\sigma\|_{\mathrm{op}}\int_{|z|\geq 1}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z=\|\sigma\|_{\mathrm{op}}\frac{A(d,\alpha)\omega_{d-1}}{\alpha-1}.$
-A straightforward calculation shows for all $\xi,\eta\in{\mathds{R}^{d}}$,
-$\left\langle\nabla^{2}f(\xi),\eta\eta^{\top}\right\rangle_{\mathrm{HS}}$ $=(1+|\xi|^{2})^{-3/2}\sum_{i,j=1}^{d}\left[(1+|\xi|^{2})\delta_{ij}-\xi_{i}\xi_{j}\right]\eta_{i}\eta_{j}$
-$=(1+|\xi|^{2})^{-3/2}\left(\sum_{i=1}^{d}\left[1+|\xi|^{2}-\xi_{i}^{2}\right]\eta_{i}^{2}-\sum_{i=1}^{d}\xi_{i}\eta_{i}\sum_{j\neq i}\xi_{j}\eta_{j}\right)$
-$\leq(1+|\xi|^{2})^{-3/2}\left((1+|\xi|^{2})\sum_{i=1}^{d}\eta_{i}^{2}+\left(\sum_{i=1}^{d}|\xi_{i}\eta_{i}|\right)^{2}\right)$
-$\leq(1+|\xi|^{2})^{-3/2}\left((1+|\xi|^{2})|\eta|^{2}+|\xi|^{2}|\eta|^{2}\right)$
-$\leq 2(1+|\xi|^{2})^{-1/2}|\eta|^{2}$
-$\leq 2|\eta|^{2}.$
-Thus, we get for all $z\in{\mathds{R}^{d}}$,
-$f(X_{t}^{x}+\sigma z)-f(X_{t}^{x})-\langle\nabla f(X_{t}^{x}),\sigma z\rangle$ $=\int_{0}^{1}\left\langle\nabla^{2}f(X_{t}^{x}+r\sigma z),(\sigma z)(\sigma z)^{\top}\right\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r$
-$\leq 2|\sigma z|^{2}\int_{0}^{1}(1-r)\,\mathrm{d}r$
-$=|\sigma z|^{2}.$
-This implies that
-$\mathsf{J}_{3}\leq\|\sigma\|_{\mathrm{op}}^{2}\int_{|z|<1}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z=\|\sigma\|_{\mathrm{op}}^{2}\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}.$
-Now we combine the estimates for the terms $\mathsf{J}_{i}$, $i=1,2,3$, and apply (1.6).
-Just as in the proof of Lemma 5.1 (where the second moment is now replaced by $\mathds{E}f(X_{t}^{x})$) we conclude that for all $\alpha\in(1,2)$,
-$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathds{E}f(X_{t}^{x})$ $\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+C_{2}+\|\sigma\|_{\mathrm{op}}\frac{A(d,\alpha)\omega_{d-1}}{\alpha-1}+\|\sigma\|_{\mathrm{op}}^{2}\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
-$\leq-\frac{\theta_{0}}{2}f(X_{t}^{x})+\frac{Cd}{\alpha-1}.$
-Gronwall’s inequality finally yields
-$\mathds{E}f(X_{t}^{x})\leq\mathrm{e}^{-\theta_{0}t/2}f(x)+\frac{2}{\theta_{0}}\,\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right).\qed$
-Proposition 5.3.
-Assume (H1). Then
-$\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)\leq C\sqrt{d}\quad\text{and}\quad\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)\leq\frac{Cd}{\alpha-1}\text{\ \ for all $\alpha\in(1,2)$}.$
-Proof.
-By the classical ergodic theory for Markov processes, see e.g. [11], we have for all $x,y\in{\mathds{R}^{d}}$,
-$\lim_{t\to\infty}W_{1}\left(\mathrm{law}(X_{t}^{x}),\mu_{\alpha}\right)=\lim_{t\to\infty}W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)=0,$
-where $W_{1}$ is the Wasserstein-$1$ distance, which is for two probability measures $\mu,\nu$ given by
-$W_{1}(\mu,\nu)=\sup_{f\in\mathrm{Lip}(1)}|\mu(f)-\nu(f)|.$
-By the triangle inequality and Lemma 5.1, we obtain for all $y\in{\mathds{R}^{d}}$ and $t\geq 0$,
-$\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)$ $\leq\left|\mathds{E}\left[|Y_{t}^{y}|\right]-\int_{\mathds{R}^{d}}|z|\,\mu_{2}(\mathrm{d}z)\right|+\mathds{E}\left[|Y_{t}^{y}|\right]$
-$\leq W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)+\sqrt{\mathds{E}\left[|Y_{t}^{y}|^{2}\right]}$
-$\leq W_{1}\left(\mathrm{law}(Y_{t}^{y}),\mu_{2}\right)+\mathrm{e}^{-\theta_{0}t/2}|y|+C\sqrt{d\left(1-\mathrm{e}^{-\theta_{0}t}\right)}.$
-Letting $t\to\infty$ we get the first assertion. Similarly, using Lemma 5.2, we have for all $x\in{\mathds{R}^{d}}$, $t\geq 0$ and $\alpha\in(1,2)$,
-$\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)$ $\leq\left|\mathds{E}\left[|X_{t}^{x}|\right]-\int_{\mathds{R}^{d}}|z|\,\mu_{\alpha}(\mathrm{d}z)\right|+\mathds{E}\left[|X_{t}^{x}|\right]$
-$\leq W_{1}\left(\mathrm{law}(X_{t}^{x}),\mu_{\alpha}\right)+\mathrm{e}^{-\theta_{0}t/2}\left(1+|x|^{2}\right)^{1/2}+\frac{Cd}{\alpha-1}\left(1-\mathrm{e}^{-\theta_{0}t/2}\right).$
-Letting $t\to\infty$ finishes the proof. ∎
-6. Proof of Theorem 1.2
-As before, $C,C_{1},C_{2}$ denote constants whose value may change from line to line.
-They depend only on the parameters $\theta_{0},\theta_{1},\theta_{2},K$ from (H1) and (H2).
-If we want to stress further dependences, we write $C(\kappa),C_{1}(\kappa),C_{2}(\kappa)$ etc.
-6.1. Gradient estimates for stable semigroup $P_{t}$
-Proposition 6.1.
-Assume (H1) and (1.3), and let $\sigma=I_{d\times d}$ and $\alpha_{0}\in(1,2)$ be an arbitrary fixed number. For all $\alpha\in[\alpha_{0},2)$, $h\in C({\mathds{R}^{d}},\real)$ and $t>0$, we have
-(6.1) $\|\nabla P_{t}h\|_{\infty}\leq C_{1}(\alpha_{0})\left\{t^{-1/\alpha}\vee 1\right\}\mathrm{e}^{-C_{2}(\alpha_{0})t}\|h\|_{\infty}.$
-Proof.
-Let $S_{t}$ be an $\frac{\alpha}{2}$-stable subordinator with Laplace transform
-$\mathds{E}\left[\mathrm{e}^{-rS_{t}}\right]=\mathrm{e}^{-2^{-1}t(2r)^{\alpha/2}},\quad r>0,\,t\geq 0.$
-By [3, Corollary 1.2] or [17, Theorem 2.1  (2)], we have for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(0,2)$ and $t>0$,
-$\|\nabla P_{t}h\|_{\infty}^{2}\leq\mathds{E}\left[\left(\int_{0}^{t}\mathrm{e}^{-2\theta_{1}r}\,\mathrm{d}S_{r}\right)^{-1}\right]\|h\|_{\infty}\leq\mathrm{e}^{2\theta_{1}t}\mathds{E}\left[S_{t}^{-1}\right]\|h\|_{\infty}.$
-Combining this with the moment identity (2.11), we obtain that for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(0,2)$ and $t>0$,
-[…]
-This means that for any $h\in C_{b}({\mathds{R}^{d}},\real)$, $\alpha\in(1,2)$ and $t\in(0,1]$,
-(6.2) $\|\nabla P_{t}h\|_{\infty}\leq Ct^{-1/\alpha}\|h\|_{\infty}.$
-Thus, (6.1) holds for $t\in(0,1]$ and it remains to prove (6.1) for $t>1$. By Lemma 6.2 and Remark 6.3 below, we have for all $\alpha\in[\alpha_{0},2)$, $x,y\in{}^{d}$ and $t>0$,
-$\sup_{f\in\mathrm{Lip}(1)}|P_{t}f(x)-P_{t}f(y)|\leq C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})t}|x-y|,$
-which means that for any $f\in\mathrm{Lip}(1)$,
-$\|\nabla P_{t}f\|_{\infty}\leq C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})t}\quad\text{for all $t>0$}.$
-Since (6.2) implies that $P_{1}h$ is a Lipschitz function with Lipschitz constant $C\|h\|_{\infty}$, we have
-$\|\nabla P_{t}h\|_{\infty}=\|\nabla P_{t-1}\{P_{1}h\}\|_{\infty}\leq CC_{1}(\alpha_{0})\|h\|_{\infty}\mathrm{e}^{-C_{2}(\alpha_{0})(t-1)},$
-which implies (6.1) for $t>1$. ∎
-By a careful calculation following the proof of [19, Theorem 1.2 (i)], we get the following lemma. To save space, we omit the detailed proof here.
-Lemma 6.2.
-Consider the SDE (1.1) with $\sigma=I_{d\times d}$ and assume that $b:{}^{d}\to{}^{d}$ is a continuous vector field such that for any $x,y\in{}^{d}$,
-(6.3) $\langle b(x)-b(y),x-y\rangle\leq\begin{cases}K_{1}|x-y|^{2},&|x-y|\leq L_{0},\\ -K_{2}|x-y|^{2},&|x-y|>L_{0}\end{cases}$
-holds with some positive constants $K_{1},K_{2},L_{0}>0$.
-Let $\alpha_{0}\in(1,2)$ be an arbitrary fixed number.
-Then there exist constants $C_{i}(\alpha_{0},K_{1},K_{2},L_{0})>0$, $i=1,2$, such that for all $\alpha\in[\alpha_{0},2)$, $x,y\in{}^{d}$ and $t>0$,
-$W_{1}\left(\mathrm{law}(X_{t}^{x}),\mathrm{law}(X_{t}^{y})\right)\leq C_{1}(\alpha_{0},K_{1},K_{2},L_{0})\mathrm{e}^{-C_{2}(\alpha_{0},K_{1},K_{2},L_{0})t}|x-y|.$
-Remark 6.3.
-It is clear that (6.3) follows from (H1) and (1.3); in this case $K_{1}=\theta_{1}$, $K_{2}=\theta_{0}/2$ and $L_{0}=\sqrt{2K/\theta_{0}}$.
-6.2. Proof of Theorem 1.2
-Proof.
-a) By [21, Theorem 1.2 (i)] with $V(x)=1+|x|$ (see also [1, Proposition 1.5]), we know for $\alpha\in(1,2)$,
-$\sup_{g:\|g/V\|_{\infty}\leq 1}|P_{t}g(x)-\mu_{\alpha}(g)|\leq C_{1}(d,\alpha)\mathrm{e}^{-C_{2}(d,\alpha)t}V(x),$
-which immediately implies that for $h\in C_{b}({\mathds{R}^{d}},\real)$ and $\alpha\in(1,2)$,
-(6.4) $|P_{t}h(x)-\mu_{\alpha}(h)|\leq C_{1}(d,\alpha)\mathrm{e}^{-C_{2}(d,\alpha)t}\|h\|_{\infty}(1+|x|).$
-It is a well-known result that $Y_{t}$ is ergodic, see e.g. [11], and for $h\in C_{b}({\mathds{R}^{d}},\real)$,
-(6.5) $|Q_{t}h(x)-\mu_{2}(h)|\leq C_{1}(d)\mathrm{e}^{-C_{2}(d)t}\|h\|_{\infty}(1+|x|).$
-Now we can use a standard method (see e.g. [7, Proposition 6.1]) to show that
-(6.6) $f_{\alpha}(x)$ $=\int_{0}^{\infty}\left[P_{t}h(x)-\mu_{\alpha}(h)\right]\mathrm{d}t,$
-(6.7) $f_{2}(x)$ $=\int_{0}^{\infty}\left[Q_{t}h(x)-\mu_{2}(h)\right]\mathrm{d}t$
-solve (1.8) for $\alpha\in(1,2]$ and $h\in C_{b}({\mathds{R}^{d}},\real)$. It follows from (6.6) and (6.4) that for all $\alpha\in(1,2)$,
-$|f_{\alpha}(x)|\leq\int_{0}^{\infty}\left|P_{t}h(x)-\mu_{\alpha}(h)\right|\mathrm{d}t$ $\leq C_{1}(d,\alpha)\|h\|_{\infty}(1+|x|)\int_{0}^{\infty}\mathrm{e}^{-C_{2}(d,\alpha)t}\,\mathrm{d}t$
-$=\frac{C_{1}(d,\alpha)}{C_{2}(d,\alpha)}\|h\|_{\infty}(1+|x|),$
-hence $f_{\alpha}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$. Similarly, we deduce from (6.7) and (6.5) that $f_{2}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$.
-b) Let $h\in C_{b}({\mathds{R}^{d}},\real)$. Without loss of generality, we may assume that $\|h\|_{\infty}\leq 1$, otherwise we replace $h$ by $h/\|h\|_{\infty}$. From (6.6) and (6.7) we have
-$|f_{\alpha}(x)-f_{2}(x)|$ $\leq\int_{0}^{\infty}\left|P_{t}h(x)-Q_{t}h(x)-\mu_{\alpha}(h)+\mu_{2}(h)\right|\mathrm{d}t$
-$\leq\left|\mu_{\alpha}(h)-\mu_{2}(h)\right|A+\int_{0}^{A}\left|Q_{t}h(x)-P_{t}h(x)\right|\mathrm{d}t$
-$\quad\mbox{}+\int_{A}^{\infty}\left|Q_{t}h(x)-\mu_{2}(h)\right|\mathrm{d}t+\int_{A}^{\infty}\left|P_{t}h(x)-\mu_{\alpha}(h)\right|\mathrm{d}t$
-$=:\mathsf{I}_{0}+\mathsf{I}_{1}(x)+\mathsf{I}_{2}(x)+\mathsf{I}_{3}(x).$
-The number $A\geq 1$ will be chosen later on. We estimate $\mathsf{I}_{0},\dots,\mathsf{I}_{4}(x)$ separately.
-From (1.7) we know that for all $\alpha\in[\alpha_{0},2)$,
-$\mathsf{I}_{0}\leq\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}A\leq C(\alpha_{0})d^{3/2}(2-\alpha)A.$
-For $\mathsf{I}_{1}(x)$ we can use Theorem 1.1 and find for all $\alpha\in[\alpha_{0},2)$ and $t>0$,
-$\left|Q_{t}h(x)-P_{t}h(x)\right|$ $\leq\|\mathrm{law}(X^{x}_{t})-\mathrm{law}(Y^{x}_{t})\|_{\mathrm{TV}}$
-$\leq C(\alpha_{0})d^{3/2}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
-Since for all $\alpha\in[\alpha_{0},2)$,
-$\int_{0}^{A}\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\mathrm{d}t\leq\int_{0}^{1}\left(1+t^{-\frac{1}{\alpha_{0}}+\frac{1}{2}}\log\tfrac{1}{t}\right)\mathrm{d}t+\int_{1}^{A}\mathrm{d}t\leq C(\alpha_{0})A,$
-we get
-$\mathsf{I}_{1}(x)\leq C(\alpha_{0})d^{3/2}(2-\alpha)A,\quad x\in{\mathds{R}^{d}}.$
-Because $\mu_{2}$ is an invariant measure of $Q_{t}$, we have for $\mathsf{I}_{2}(x)$
-$\int_{A}^{\infty}\left|Q_{t}h(x)-\mu_{2}(h)\right|\mathrm{d}t=\int_{A}^{\infty}\left|\int_{{\mathds{R}^{d}}}\left[Q_{t}h(x)-Q_{t}h(y)\right]\mu_{2}(\mathrm{d}y)\right|\mathrm{d}t.$
-For all $t\geq 1$ we have
-$\left|Q_{t}h(x)-Q_{t}h(y)\right|$ $=\left|\int_{0}^{1}\langle\nabla Q_{t}h(rx+(1-r)y),x-y\rangle\,\mathrm{d}r\right|$
-$\leq\|\nabla Q_{t}h\|_{\infty}|x-y|$
-$\leq C_{1}\mathrm{e}^{-C_{2}t}(|x|+|y|),$
-where the last inequality is due to (3.1). Hence,
-$\mathsf{I}_{2}(x)$ $\leq\int_{A}^{\infty}\int_{\mathds{R}^{d}}|Q_{t}h(x)-Q_{t}h(y)|\,\mu_{2}(\mathrm{d}y)\,\mathrm{d}t$
-$\leq C_{1}\int_{\mathds{R}^{d}}(|x|+|y|)\,\mu_{2}(\mathrm{d}y)\int_{A}^{\infty}\mathrm{e}^{-C_{2}t}\,\mathrm{d}t$
-$=\frac{C_{1}}{C_{2}}\,\mathrm{e}^{-C_{2}A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{2}(\mathrm{d}y)\right].$
-Replacing (3.1) in the above argument with (6.1) yields for all $\alpha\in[\alpha_{0},2)$ the estimate for $\mathsf{I}_{3}(x)$:
-$\mathsf{I}_{3}(x)$ $=\int_{A}^{\infty}\left|\int_{\mathds{R}^{d}}[P_{t}h(x)-P_{t}h(y)]\,\mu_{\alpha}(\mathrm{d}y)\right|\mathrm{d}t$
-$\leq\int_{A}^{\infty}\int_{\mathds{R}^{d}}|P_{t}h(x)-P_{t}h(y)|\,\mu_{\alpha}(\mathrm{d}y)\,\mathrm{d}t$
-$\leq C_{1}(\alpha_{0})\int_{\mathds{R}^{d}}(|x|+|y|)\,\mu_{\alpha}(\mathrm{d}y)\int_{A}^{\infty}\mathrm{e}^{-C_{2}(\alpha_{0})t}\,\mathrm{d}t$
-$=\frac{C_{1}(\alpha_{0})}{C_{2}(\alpha_{0})}\,\mathrm{e}^{-C_{2}(\alpha_{0})A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{\alpha}(\mathrm{d}y)\right].$
-Altogether we have shown that for all $\alpha\in[\alpha_{0},2)$,
-$|f_{\alpha}(x)-f_{2}(x)|$
-$\quad\leq C(\alpha_{0})d^{3/2}(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}\left[|x|+\int_{\mathds{R}^{d}}|y|\,\mu_{2}(\mathrm{d}y)+\int_{\mathds{R}^{d}}|y|\,\mu_{\alpha}(\mathrm{d}y)\right]$
-$\quad\leq C(\alpha_{0})d^{3/2}(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}d(1+|x|)$
-$\quad\leq d^{3/2}\left[C(\alpha_{0})(2-\alpha)A+C_{1}(\alpha_{0})\mathrm{e}^{-C_{2}(\alpha_{0})A}\right](1+|x|),$
-where the second inequality is due to Proposition 5.3. Taking $A=1\vee[|\log(2-\alpha)|/C_{2}(\alpha_{0})]$, we get for all $\alpha\in[\alpha_{0},2)$ and $x\in{\mathds{R}^{d}}$,
-$|f_{\alpha}(x)-f_{2}(x)|\leq C(\alpha_{0})d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha}\cdot(1+|x|).\qed$
-7. Exact rate for the Ornstein–Uhlenbeck process on
+Proof of Corollary 1.2.
+It follows from [8, Lemma 2.1] with $\psi(r)=r^{p}$ that for any $p\in(0,1)$ and probability measures $\mu,\nu$,
+$\|\mu-\nu\|_{W_{p}}$ $\leq\inf_{t>0}\left\{t^{p/2}\sqrt{d}\,\|\mu-\nu\|_{\mathrm{TV}}+t^{(p-1)/2}d\,\|\mu-\nu\|_{W_{1}}\right\}$
+$\leq d^{(p+1)/2}\left\{\|\mu-\nu\|_{\mathrm{TV}}+\|\mu-\nu\|_{W_{1}}\right\}.$
+If (H1) and (H2) hold, we know from [5, Theorem 1.1] that for any $\alpha\in(1,2)$,
+$\|\mu_{\alpha}-\mu_{2}\|_{W_{1}}\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
+Combining these two estimates with (1.6), we obtain the desired estimate. ∎
+5. Optimality for the Ornstein–Uhlenbeck process on
 In this section, we assume that $\mu_{\alpha}$ is the invariant measure of the Ornstein–Uhlenbeck process on , which solves the following SDE:
@@
 where $B_{t}$ is a standard Brownian motion on with $\mathds{E}\left[\mathrm{e}^{\mathop{\mathrm{i}}\xi B_{t}}\right]=\mathrm{e}^{-t|\xi|^{2}/2}$.
-It is easy to see that (H1) holds with $\theta_{0}=1$ and $K=0$, while (H2) holds with $\theta_{1}=1$ and $\theta_{2}=0$.
+It is easy to see that (H1) holds with $\sigma=1$, while (H2) holds with $\theta_{0}=1$, $K=0$, $\theta_{1}=1$ and $\theta_{2}=0$.
 Proposition 5.1.
@@
 $\liminf_{\alpha\uparrow 2}\frac{\|\mu_{2}-\mu_{\alpha}\|_{\mathrm{TV}}}{2-\alpha}\geq\lim_{\alpha\uparrow 2}\frac{\mathrm{e}^{-1/4}-\mathrm{e}^{-1/(2\alpha)}}{2-\alpha}=\frac{1}{8}\,\mathrm{e}^{-1/4}.\qed$
+6. Appendix
+The following exponential contraction result is due to Eberle [6, Corollary 2], but in concrete situations it is very hard to verify the rather general conditions of that paper without examining the proofs.
+Therefore, and for the convenience of our readers, we decided to give a short proof of our own, based on Eberle’s ideas, which is very easy to apply in the present situation.
+Proposition 6.1.
+Consider the SDE (1.2) and assume (H1), (1.3) and (1.4). There exist constants $C_{1},C_{2}>0$ such that for all $x,y\in{\mathds{R}^{d}}$ and $t\geq 0$,
+$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{W_{1}}\leq C_{1}\mathrm{e}^{-C_{2}t}|x-y|.$
+We begin with a few preparations. Set $R_{0}:=(2K/\theta_{0})^{1/2}$. It follows from (1.3) and (1.4) that for all $x,y\in{\mathds{R}^{d}}$,
+(6.1) $\langle x-y,b(x)-b(y)\rangle\leq\phi(|x-y|)|x-y|,$
+where
+$\phi(r):=\theta_{1}r\mathds{1}_{[0,R_{0}]}(r)-\frac{1}{2}\theta_{0}r\mathds{1}_{(R_{0},\infty)}(r),\quad r\geq 0.$
+Without loss of generality, we may assume that the matrix $\sigma\in\real^{d\times d}$ is symmetric and positive definite; otherwise, we could replace $\sigma$ by $(\sigma\sigma^{\top})^{1/2}$.
+Denote by $\lambda_{\min}$ the smallest eigenvalue of $\sigma$.
+For $r\geq 0$, let
+$f(r):=\int_{0}^{r}\exp\left[-\frac{\theta_{0}+2\theta_{1}}{8\lambda_{\min}}(s\wedge R_{0})^{2}\right]\mathrm{d}s.$
+It is easy to verify that $f$ is piecewise of class $C^{2}$, $f^{\prime}>0$ is continuous and decreasing on $0,\infty)$, and
+(6.2) $f^{\prime}(R_{0})r\leq f(r)\leq r,\quad r\geq 0.$
+Lemma 6.2.
+For all $r\in[0,\infty)\setminus\{R_{0}\}$,
+$f^{\prime}(r)\phi(r)+2\lambda_{\min}f^{\prime\prime}(r)\leq-\frac{\theta_{0}}{2}f^{\prime}(R_{0})f(r).$
+Proof.
+If $r\in[0,R_{0})$, we get by the definition of $f$, the monotonicity of $f^{\prime}$ and the second inequality in (6.2),
+$f^{\prime}(r)\phi(r)+2\lambda_{\min}f^{\prime\prime}(r)$ $=\theta_{1}rf^{\prime}(r)+2\lambda_{\min}f^{\prime\prime}(r)$
+$=-\frac{\theta_{0}}{2}rf^{\prime}(r)\leq-\frac{\theta_{0}}{2}f^{\prime}(R_{0})r\leq-\frac{\theta_{0}}{2}f^{\prime}(R_{0})f(r).$
+If $r>R_{0}$, then $f^{\prime}(r)=f^{\prime}(R_{0})$ and $f^{\prime\prime}(r)=0$, and we get
+(6.3) $f^{\prime}(r)\phi(r)+2\lambda_{\min}f^{\prime\prime}(r)=f^{\prime}(R_{0})\cdot\left(-\frac{\theta_{0}}{2}r\right)\leq-\frac{\theta_{0}}{2}f^{\prime}(R_{0})f(r).$
+Proof of Proposition 6.1.
+Fix $x,y\in{\mathds{R}^{d}}$ and assume that $x\neq y$. Consider the following SDE with initial value $\hat{Y}_{0}=y$
+$\left\{\begin{aligned} \mathrm{d}\hat{Y}_{t}&=b(\hat{Y}_{t})\,\mathrm{d}t+\sigma\left(I_{d\times d}-2e_{t}e_{t}^{\top}\right)\mathrm{d}B_{t},&&t<\tau,\\ \hat{Y}_{t}&=Y_{t}^{x},&&t\geq\tau,\end{aligned}\right$.
+where $I_{d\times d}\in\real^{d\times d}$ is the identity matrix, $e_{t}:=\frac{\sigma^{-1}(Y_{t}^{x}-\hat{Y}_{t})}{|\sigma^{-1}(Y_{t}^{x}-\hat{Y}_{t})|}$ and $\tau:=\inf\left\{t\geq 0\,:\,\hat{Y}_{t}=Y_{t}^{x}\right\}$ is the coupling time.
+Since the process $(I_{d\times d}-2e_{t}e_{t}^{\top})_{t\geq 0}$ takes values in the orthogonal $d\times d$-matrices, we know from Lévy’s criterion that
+$\int_{0}^{t}\left(I_{d\times d}-2e_{s}e_{s}^{\top}\mathds{1}_{\{s<\tau\}}\right)\mathrm{d}B_{s},\quad t\geq 0$
+is a standard Brownian motion in $\mathds{R}^{d}$}.
+Therefore, $\mathrm{law}(\hat{Y}_{t})=\mathrm{law}(Y_{t}^{y})$ for all $t\geq 0$.
+Set $Z_{t}:=Y_{t}^{x}-\hat{Y}_{t}$ and note that $(W_{t})_{t\geq 0}$, $W_{t}:=\int_{0}^{t}e_{s}^{\top}\,\mathrm{d}B_{s}$, is a standard Brownian motion in .
+By the Itô-Meyer formula, we get for all $t<\tau$
+$\mathrm{d}|Z_{t}|=\frac{1}{|Z_{t}|}\left\langle Z_{t},b(Y_{t}^{x})-b(\hat{Y}_{t})\right\rangle\mathrm{d}t+\frac{2|Z_{t}|}{|\sigma^{-1}Z_{t}|}\,\mathrm{d}W_{t}$
+(note that the local-time part vanishes as $t<\tau$, i.e.
+$Z_{t}\neq 0$).
+Since $f$ is piecewise $C^{2}$ and $f^{\prime}$ is continuous, we can use Itô’s formula and the expression for $\mathrm{d}|Z_{t}|$ to get for all $t<\tau$,
+$\mathrm{d}f(|Z_{t}|)=\frac{f^{\prime}(|Z_{t}|)}{|Z_{t}|}\left\langle Z_{t},b(Y_{t}^{x})-b(\hat{Y}_{t})\right\rangle\mathrm{d}t+f^{\prime\prime}(|Z_{t}|)\frac{2|Z_{t}|^{2}}{|\sigma^{-1}Z_{t}|^{2}}\,\mathrm{d}t+f^{\prime}(|Z_{t}|)\frac{2|Z_{t}|}{|\sigma^{-1}Z_{t}|}\,\mathrm{d}W_{t}.$
+Since $f^{\prime\prime}\leq 0$ on $0,\infty)\setminus\{R_{0}\}$ and $|\sigma^{-1}x|\leq|x|/\lambda_{\min}$ for $x\in{\mathds{R}^{d}}$, it follows from (6.1) and Lemma 6.2 that
+$\mathrm{d}f(|Z_{t}|)$ $\leq\left[f^{\prime}(|Z_{t}|)\phi(|Z_{t}|)+2\lambda_{\min}f^{\prime\prime}(|Z_{t}|)\right]\mathrm{d}t+f^{\prime}(|Z_{t}|)\frac{2|Z_{t}|}{|\sigma^{-1}Z_{t}|}\,\mathrm{d}W_{t}$
+$\leq-Cf(|Z_{t}|)\,\mathrm{d}t+f^{\prime}(|Z_{t}|)\frac{2|Z_{t}|}{|\sigma^{-1}Z_{t}|}\,\mathrm{d}W_{t},$
+where $C=\frac{1}{2}\theta_{0}f^{\prime}(R_{0})$. This implies for $t<\tau$ that
+$\mathrm{d}\left[\mathrm{e}^{Ct}f(|Z_{t}|)\right]\leq\mathrm{e}^{Ct}f^{\prime}(|Z_{t}|)\frac{2|Z_{t}|}{|\sigma^{-1}Z_{t}|}\,\mathrm{d}W_{t}.$
+Using the finite stopping times $\tau_{n}:=\inf\left\{t\geq 0\,:\,|Z_{t}|\notin[1/n,n]\right\}\leq\tau$ and optional stopping, we see that the right-hand side of the above inequality is a mean-zero martingale.
+Hence,
+$\mathds{E}\left[\mathrm{e}^{C(t\wedge\tau_{n})}f(|Z_{t\wedge\tau_{n}}|)\right]\leq f(|Z_{0}|)=f(|x-y|),\quad t\geq 0.$
+Since $\tau_{n}\uparrow\tau$ as $n\uparrow\infty$, we can apply Fatou’s lemma and get
+$\mathds{E}\left[\mathrm{e}^{C(t\wedge\tau)}f(|Z_{t\wedge\tau}|)\right]\leq f(|x-y|).$
+Noting that $Z_{t}=0$ for $t\geq\tau$ and $f(0)=0$, we conclude that
+$\mathrm{e}^{Ct}\mathds{E}f(|Z_{t}|)\leq f(|x-y|).$
+Because of (6.2), we see that for all $t\geq 0$,
+$f^{\prime}(R_{0})\mathds{E}|Y_{t}^{x}-\hat{Y}_{t}|\leq\mathds{E}f(|Z_{t}|)\leq\mathrm{e}^{-Ct}f(|x-y|)\leq\mathrm{e}^{-Ct}|x-y|.$
+Since, by construction, $\mathrm{law}(\hat{Y}_{t})=\mathrm{law}(Y_{t}^{y})$, the proof is finished. ∎
 Acknowledgement.
@@
 Schilling is supported by the 6G-Life project (BMBF Grant No.
-16KISK001K) and the ScaDS.AI centre.
+16KISK001K) and TU Dresden’s ScaDS.AI initiative.
 L.